### v1

# Space Group Informed Transformer for Crystalline Materials Generation

Abstract
We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials. The space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials. Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and locations of symmetry-inequivalent atoms in the unit cell. Our results demonstrate that CrystalFormer matches state-of-the-art performance on standard benchmarks for both validity, novelty, and stability of the generated crystalline materials. Our analysis also shows that CrystalFormer ingests sensible solid-state chemistry information from data for generative modeling. The CrystalFormer unifies symmetry-based structure search and generative pre-training in the realm of crystalline materials. The simplicity, generality, and flexibility of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials modeling and discovery.
I Introduction
Machine learning methods are playing an increasingly important role in material discovery, complementing conventional computational approaches [1, 2]. Generative machine learning, in particular, has been a promising step for matter inverse design [3, 4] which goes beyond machine learning accelerated structure search [5] and property screening [6]. Generative models learn the underlying distribution of training data and generated new samples from the learned distribution. In addition, the generation process can also be controlled by prompts such as desired material properties or experiment observations. Amazing programming abilities of generative models have been demonstrated in large language model [7], text-to-image generation [8, 9], and protein design [10].
It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery. A generative pretrained foundation model for crystalline materials is a key step towards such lofty goal. However, despite of intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains. Simply scaling the compute and model size of the current crystal generative model may not be enough because the amount of high-quality data for crystalline materials is much less than compared to language and image domains. Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [12, 23, 24, 25, 26].
| world | With space group symmetry |
The space group symmetry is arguably the most important inductive bias in the modeling of crystalline materials, which is the joint outcome of the rotational and translational symmetry in space. There are in total 230 space groups [29] for three-dimensional crystal structures. Nature exhibits a preference for symmetric crystal structures, a tendency that may be attributed to the symmetry inherent in the interatomic interactions, which, in turn, are governed by the fundamental forces acting between elementary particles. As a result, the appearance of crystalline materials in the the first and the least symmetric space group is rare [30], with many instances potentially being misclassified. [31].
Space group symmetry imposes significant constraints on a crystal. At its core, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the dimensions and angles of the crystal’s unit cell. Moreover, the symmetry operations associated with a given space group ensure that identical atoms are consistently mapped onto each other across the crystal, preserving its structural integrity. This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the total number of atoms in each unit cell. A key concept in understanding these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group. These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes. For example, Fig. 1(a) shows the Wyckoff positions for the space group (No. 167). The Wyckoff positions are labeled by letters in the alphabet, starting from special points in the bottom to general positions in the top. The multiplicity counts the number of equivalent positions connected by the space group symmetry operations. All of them should be occupied by the same type of atoms to uphold the space group symmetry. For example, the top row of the table in Fig. 1(a) shows a general position that can be mapped to 36 positions under the symmetry operations of the space group.
Nature tends to place atoms in those special Wyckoff positions at the bottom of the table. For example, we highlight the occupied Wyckoff positions of calcite (CaCO) crystal in Fig. 1, associated with the space group. One sees that the Wyckoff letter ’6a’ and ’6b’ deterministicly define the locations of the carbon and calcium atoms within the unit cell. In addition, it follows that , and as the space group belongs to the trigonal crystal system. Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO structure: the coordinate of oxygen atom and the lattice constants and . All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
The prominent discrete and sequential features illustrated in Figure 1 are ubiquitous in crystalline materials. The Wyckoff positions not only specify possible locations of atoms in the unit cell, their associated multiplicities also put strong constraints on the number of atoms. Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials. Failing to exploit this information in generative modeling not only renders learning inefficient, it also severely impairs the generalization ability of the model. For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16].
In this paper, we introduce CrystalFormer, an autoregressive transformer for generative modeling of crystalline materials. CrystalFormer models the joint probability distribution of Wyckoff positions, chemical species, and lattice parameters of crystals with a given space group. By treating the Wyckoff positions as the first class citizen in the model, CrystalFormer seamlessly integrates the space group symmetry into crystal probabilistic modeling. The space group-informed transformer exploits the fundamental inductive bias of crystals to greatly constrain and simplify the generation process. As shown in Table 1, explicit modeling of the Wyckoff positions greatly reduces the space of crystalline materials. We benchmark the CrystalFormer model on standard crystal generation tasks and show it exhibits great efficiency and generalization ability compared to existing models. Finally, we discuss its potential applications in novel materials discovery. We have released the codes and trained model at [32].
II CrystalFormer
To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms. Wyckoff letters follows the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group. Later letters in the alphabet indicate more general positions with reduced site symmetries. Note that the information of the space group number and Wyckoff letter fully determine the multiplicities. In cases where the atom positions are fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g. the -coordinate of the oxygen atoms in the CaCO example shown in Fig. 1, To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially. The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry regions with larger multiplicities.
With these considerations, we define a crystal data as . Here are Wyckoff letters and are chemical species. Here, stands for the number of symmetrically inequivalent atoms in the conventional unit cell. For example, as shown in Fig. 1(b) one has for CaCO. Explicitly including the Wyckoff letter in the generative modeling is the key of the present work. Next, are the fractional coordinates of symmetrically inequivalent atoms. Lastly, denotes the lattice parameters of the conventional unit cell of the material.
The central quantity to focus on is the conditional probability of a crystal given the space group number : . Since the space group is a fundamental characterization for crystalline materials, is a key control variable that greatly simplifies the distribution over the entire crystal materials space. In practical applications, the space group can be considered separately or even predicted [33, 34].
We express the space group conditioned probability distribution of crystals as an autoregressive product of conditional probabilities
At first sight, it may appear unnatural to employ an autoregressive model for crystals since there is no obvious order for atoms in the unit cell. However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters. During sampling, the autoregressive model places the atoms one by one into the unit cell, starting from the special position with high symmetry to the general position with the lowest symmetry, see Fig. 1(c).
We model the conditional probability of the Wyckoff letters and chemical species as categorical distributions. A mixture of von-Mises distribution is employed to model the conditional probability of the factional coordinates . For Wyckoff positions with multiplicities greater than one, we only consider the first of fractional coordinates that appear in the international tables for crystallography [35]. Lastly, we model the conditioned distribution of lattice parameters as a Gaussian mixture model.
We build CrystalFormer using the vanilla decoder-only transformer [36] to model the probability distribution of crystalline materials Eq. (1). The transformer-based autoregressive model is a favorable choice to capture these prominent discrete characteristics of crystals. We feed atomistic features into an autoregressive transformer. The space group number is the first input to the transformer. The remaining inputs alternate between the Wyckoff letter, chemical species, and fractional coordinates of each atom. One can go through the table of Fig. 1(b) in a raster order to collect these atomistic features. We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer. In particular, we concatenate the vector embedding of to all other inputs since it is the key control variable for the crystal generation. In addition, we have also provided the multiplicity of each Wyckoff position as an additional feature. The multiplicity can be easily inferred from the space group and the Wyckoff letters. We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [37, 13] We pad the atom sequence up to a maximum length and take the shifted output as parameters of the conditional probability distribution Eq. (1), see Fig. 1(b). At the location of the first padding atom, we predict the lattice parameters.
We implement a number of constraints in the model to further reduce the phase space of the model. First, the Wyckoff letters should be valid for the given space group. For example, for the space group number 167 the Wyckoff letters go from ’a’ up to ’f’. Second, we require that the Wyckoff letters follow alphabetical order in the sequence 111We follow the convention that capital letters appear after lower case letters. This handles the edge case of the No. 47 space group whose Wyckoff positions reach ’A’. This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’.. Lastly, the Wyckoff positions with no continuous degree of freedom (such as ’a’, ’b’, and ’d’ positions in the space group 167) can only be occupied once. Those constraints are implemented by tuning the logit biases of Wyckoff letters [39, 40].
The design of CrystalFormer focuses mostly on the space group symmetries since we believe they are the most crucial ones for crystalline materials. This design decision significantly impacts the treatment of other symmetries. First, it is often possible to place the origin of the unit cell at the inversion center of the specified space group. The chosen origin naturally fixes the continuous translation invariance of fractional coordinates. Second, by only considering symmetry-inequivalent atoms and labeling them with Wyckoff letters, one fixes most of the permutation invariance in the representation. For those Wyckoff positions with continuous degrees of freedom, there may be multiple symmetry-inequivalent atoms with the same Wyckoff letters. We arrange these atoms according to the lexicographic order of fractional coordinates [41] in the sequence. Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry. Lastly, the periodicity of the fractional coordinates is respected since they are treated as variables under the von-Mises distribution.
The CrystalFormer is trained by minimizing the negative log-likelihood over training dataset
Writing out according to Eq. (1), the objective function contains the negative log-likelihood of discrete variables such as Wyckoff letters and chemical species , as well as continuous variables such as fractional coordinates and the lattice parameters . In the objective function, for continuous variables we consider only active ones that are not fixed by the space groups and Wyckoff letters. In this way, those special fractional coordinates (e.g. , ) and lattice parameters (e.g. , ) do not contribute to the loss function.
To sample crystals from the CrystalFormer, one needs to specify a space group number and a number of possible chemical elements. The CrystalFormer samples the atoms one by one, starting from more symmetric specific positions with lower multiplicities till less symmetric general positions with larger multiplicities. We use the information of the space group and Wyckoff letter to control the sampling of fractional coordinates. By applying the symmetry projection to the sampled fractional coordinate, one ensures the generated fractional coordinates are compatible with the Wyckoff positions. One can also mask out the logits of chemical species so that only a number of selected elements will be sampled. The number of symmetrically inequalivent atoms may fluctuate in the sampling procedure. Once one has sampled a padding atom, the model predicts the lattice parameters under the space group constraint. Moreover, we introduce a temperature parameter in the sample distribution . With we will draw samples from a sharper distribution, while gives more diversity in the generated samples.
III Results and Evaluations
We train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell. The training dataset contains 27136 crystal structures. The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category. On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalance distribution of crystals over space groups in Nature [30]. In fact, there is no training data in 61 out of 230 space groups as shown in Fig. 6. Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature.
Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters. Next, we will select the model checkpoint with the lowest total validation loss and examine its learned features. Then, we will use the model to generate 1000 crystal samples for each of the 230 space groups and evaluate their validity, stability, and novelty. These results highlight the benefits of built-in space group symmetries in the model and point to possible future applications of the CrystalFormer model.
III.1 Learned features
Figure 3 visualizes the cosine similarity of the learned vector embedding of the chemical species. Red colors in the figure indicate similar chemical species identified by the model. One clearly sees the periodicity shows up as off-diagonal stripes. Moreover, there are visible clusters for Lanthanide elements (La-Lu). The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc). Being able to learn chemical similarities from data [42, 43, 44, 45, 46, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures. In particular, the features shown in Fig. 3 is strikingly similar to the similarity map constructed purposely based on substitution pattern [42] which was later used for substitution-based material discovery [47].
Figure 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups. One sees a nice agreement between the atom number distribution in the test dataset and the generated samples. In addition, it appears that space group is the key latent variable that decomposes the multi-modal atom number distribution of crystals. This is understandable because the number of atoms is determined by the sum of the multiplicities of occupied Wyckoff positions. Therefore, the space group symmetry is a key control variable for the atom number distribution. Building Wyckoff positions information into the CrystalFormer model architecture removes the necessities of querying the training data to find out the number of atoms for a targeted space group [16].
Figure 5 shows heat maps of Wyckoff positions and chemical species for the space group (No. 225). First of all, one sees that most atoms occupy those special Wyckoff positions with high symmetries. Moreover, the heat map reveals interesting solid-state chemistry knowledge as it suggests where each atom tends to appear in a crystal. For example, there are vertical stripes at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials. Moreover, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities. Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials. On the other hand, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset. We believe that is due to that the hydrogen element takes only about in the training data for the space group. Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
There has been a long history of summarizing empirical or statistical chemistry rules encoded in materials data and then use them to instruct search of crystal structures [48, 49, 50, 51, 42, 52]. Our analysis shows that CrystalFormer also ingests a number of chemical intuitions in the training data for generative modelling. Since training procedure compresses the chemistry knowledge into a neural network model with tractable likelihoods, it is possible to employ CrystalFormer in the Monte Carlo or evolution strategy search of crystal structures, besides as a probabilistic generative model.
III.2 Validity and novelty of generated samples
Figure 6 illustrates the structure and compositional validity of generated samples across all 230 space groups. Following the Ref. [53], a structure meets the validity criteria if the shortest atomic distance exceeds , a lenient standard. Composition validity requires charge neutrality as computed by SMACT [54]. This is, however, an overly stringent criterion since the composition validity of the training set is only around by this measure [11]. Note that the CrystalFormer is able to generate reasonable samples even for those space groups without any training data.
| Space group | Crystal system | Training samples | Validity () | |
| Struc. | Comp. | |||
| 2 | Triclinic | 676 | 83.10 | 83.0 |
| 12 | Monoclinic | 1273 | 87.70 | 81.80 |
| 62 | Orthorhombic | 1187 | 95.50 | 87.20 |
| 139 | Tetragonal | 1233 | 97.70 | 83.40 |
| 166 | Trigonal | 1076 | 98.50 | 85.0 |
| 194 | Hexagonal | 1129 | 99.40 | 89.90 |
| 225 | Cubic | 3960 | 99.60 | 93.50 |
| 1 | Triclinic | 27136 | 91.40 | 80.20 |
| Autoregressive models | ||||
| PGSchNet [55] | 99.65 | 75.96 | ||
| LM-CH (character-level tokenization) [18] | 84.81 | 83.55 | ||
| LM-AC (atom coordinate-level tokenization) [18] | 95.81 | 88.87 | ||
| Crystal-LLM [20] | 96.5 | 86.3 | ||
| Diffusion models | ||||
| CDVAE [11] | 100.0 | 86.70 | ||
| DiffCSP [13] | 100.0 | 83.25 | ||
| DiffCSP++ [26] | 99.94 | 85.12 | ||
| UniMat-Large [15] | 97.2 | 89.4 | ||
Table 2 reports the validity of generated samples for selected space groups in each of the seven crystal systems. To ensure the numbers are representative, we chose the space group to be the one with the most training data for each crystal system. One sees that the model performs better for more symmetric space groups. This is a nice feature that complements existing crystal generative models, which mostly have difficulties in generating highly symmetric structures. As a reference, we also list the validity of the generated samples suppose one treats the crystals as if they are all in the space group (No. 1) with only translational symmetry. One sees the structure validity scores of space group improves compared to the one shown in Figure 6 due to increased training samples.
The second part of Table 2 shows reference results in the literature for the same validity test. In principle, the performance of the present model should fall back to the language model approaches [18, 20]. The remaining gap may be due to details such as the including the header line in the crystallographic information files (CIF), specific sampling strategy of language models, or the additional post-selection of samples [20]. In the table, the DiffCSP++ [26] is the only alternative model that exploits the space group symmetry in the generation process. However, different from us, the DiffCSP++ model does not predict Wycokff position as a part of generation process. Therefore, DiffCSP++ needs to search for template structure in the training set for generation, which may limit its generality. Besides works listed in Table 2 that reported validity scores in a comparable settings, Ref. [19] has conditioned the generation of CIF on the space group symbols in a language model setting. Ref. [16] considered space group conditioned crystal generation using a fine-tuned generative model with space group labels. Neither approach provides exact constraints on the space group, which could yield problematic structures for large systems and highly symmetric space groups. Ref. [23] considered generating symmetric crystals in their Wyckoff representation. However, the model does not consider fractional coordinates and lattice parameters, so it requires a subsequent computational search to completely determine the crystal structure. Sec. IV.2 contains a more in-depth discussion of the relationship of the present approach with related works.
Figure 7(a) shows the validity of generated samples as a function of sampling temperature. One clearly sees that reduced temperature increases the validity of samples at the cost of reducing the diversity [20]. Figure 7(b) shows the novelty and uniqueness evaluated on generated samples with temperature. Novelty quantifies the proportion of new structures in the generated samples that were unseen in the training dataset. Uniqueness represents the percentage of distinct, non-redundant structures among the generated samples [56]. One sees that across different temperatures the uniqueness remains high, indicating the model does not collapse to a mode that produces duplicated samples. On the other hand, the model produces close to novel material as temperature increases, which nicely demonstrates modal covering behavior of the maximum likelihood estimation training [57]. Having a model distribution broader than the span of the dataset is crucial for material discovery.
III.3 Stability of generative samples
Figure 8(a) shows the histogram of the log-likelihood of generated and test samples show nice agreement. We also visualize structures of a few generated samples which deems to be very likely, typical, and unlikely according to their log-likelihoods. Interestingly, we did not observe a clear correlation between the model likelihood and energy of the structure.
Fig. 8(b) shows the energy above hull based on energies calculated with density functional theory (DFT). One sees the distributions of unrelaxed and relaxed structures are very similar, suggesting that the generated structures are already very close to local minima of DFT calculations. For the given computational budget detailed in the Appendix B, we manage to reach convergence for 973 out of 1000 structures in the DFT relaxation. All of the structures retains the same space group symmetry after the DFT relaxation 222Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No. 221).. With its default parameters, StructureMatcher from pymatgen [60] is able to match all 973 structures before and after relaxation. The average root mean squared displacement (RMSD) normalized by computed for these matched structures is 0.0061, which confirms the generated samples are indeed very close to DFT local minima. Figure 8(b) shows a pronounced peak around eV/atom showing the model indeed generates a large fraction of candidates of metastable materials [58]. Finally, we found 21 relaxed samples are stable by the criteria and summarize them in Table S2 of appendix B. Most of the found stable materials only have atoms occupying the first few Wyckoff positions. Among these 21 structures, we found 10 are new crystalline materials that are not contained in the Materials Project database [61].
IV Discussions
IV.1 Applications
With the abilities demonstrated in Sec. III, we believe that CrystalFormer is ready to fit into the existing materials discovery workflow in the following two ways.
IV.1.1 Initialization for structure search
Crystal structure search relies heavily on utilizing space group symmetries. For example, it is a common practice for crystal structure prediction software [62, 63, 64, 65, 66] and crystal structure search [67, 68] to place atoms at Wyckoff positions to prepare symmetric initial states for subsequent optimization.
However, such initialization approach faces combinatorial difficulty as the number of chemical species and atoms in the unit cell grow. The CrystalFormer is ready to act as a drop-in replacement of random structure initialization. In this way, one bypasses the curse-of-dimensionality of exact enumeration with a data-driven probabilistic approach. Moreover, the ability of CrystalFormer to generate diverse and close to metastable structures can greatly reduce computational costs of downstream ab initio calculations.
IV.1.2 Mutation of existing structures
Mutatation of known crystals is a prominent approach to materials discovery. For example, one can employ machine machine-learned force field to relax crystal structures [69, 5, 70, 71] after element substitutions. In the lens of generative modeling, the machine learning force field can be regarded as the energy-based model or Boltzmann machines. A potential drawback of exploring materials space with an energy-based model is the slow mixing or even ergodicity issue posed by the rough landscape of the potential energy surface. In this sense, element substitutions provide a variety of initial seeds, compensating the limitation potential energy surface-based exploration.
Having an alternative measure of crystal likelihood other than the potential energy surface opens a way to employ the model likelihood as a guide for structure search. For example, with a trained CrystalFormer, one can perform Markov chain Monte Carlo (MCMC) random walk in the crystalline materials starting from an existing crystal structure. At each step of the random walk, one proposes an element substitution, atom shift, or lattice deformation and accept or reject the proposal according to the model likelihood. In the long time limit, one will sample from the model distribution. However, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
IV.2 Related works
Crystal generative models have been explored using variational autoencoder [72, 23], generative adversarial networks [73, 12], normalizing flows [74, 75, 76], diffusion models [11, 12, 13, 26, 14, 15, 16, 21, 22], GFlowNet [24, 25], and autoregressive models [77, 55, 78, 17, 18, 19, 19]. In these autoregressive models, one either use atomistic features [77, 55, 78] or use pure text tokens [17, 18, 19, 20]. However, with the introduction of specialized tokens for crystals, the boundary between the two is blurred.
The CrystalFormer is most closely related to the autoregressvie generative model originally designed for molecules [77, 55, 78]. However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of atoms in the unit cell. Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around rotational equivariance. On the other hand, compared to Refs. [17, 18, 19, 20] which treat text descriptions of crystals using autoregressive language models, we are only dealing with more concise and essential atomistic representation of crystals, which leads to smaller model size and faster sampling speed. Fast generation speed is not only a welcoming feature but also will be crucial for intelligent exploration of materials space based on combinations of probabilistic generation and post-selection, backtracking, and searching techniques [79]. More importantly, CrystalFormer guarantees space group constraints since these symmetries are baked in the model architecture rather than learned as statistical correlation from text data. In a sense, the present work employs intrinsic mathematical (as opposed to natural) language to incorporate the symmetry principle in the generative modeling of crystals.
As a side remark, the Wyckoff position features have been used in machine learning models for materials property prediction [80, 81]. Incorporating space group information in the encoder-only transformer models may also boost their property prediction performance [82, 83] as suggested by the related study [84].
V Outlook
Crystal structure prediction has long been the dream of computational material sciences [85]. CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one framework. Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials discovery.
Precisely controlling the space group in the generative model of crystalline materials is not only a highly desired feature but also greatly simplifies the task. An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups. Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16]. The transformer-based generative model is ready to be scaled up with more training data, in the same fashion as large language models [86]. Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing. When scaling up the model it will be interesting to note the possible appearance of neural scaling law [87] as it has also been showing up in other contexts of atomistic modeling [88].
The model architecture is still open to further improvement to better serve the purpose of material discovery, First of all, to better facilitate data efficiency learning and structure phase transitions-related applications, it will be useful to further exploit the group-subgroup structure [89] in the model architecture. Second, besides chemical elements, it will be useful to constrain the stoichiometry of generated materials. This may be achieved by tuning the logit bias or combining the present probabilistic generation approaches with traditional enumeration approaches [90]. Lastly, it is worth exploring to use CrystalFormer as the base distribution in the flow model and employ symmetry-persevering transportation’s to further adjust the atoms coordinates and unit cells [10, 26], which mimics a symmetry-constrained relaxation process [91].
Conditioned crystal structure generation based on material properties [78, 16] and experimental measurements [92] are both straightforward with CrystalFormer. One can either extend the space group embedding or employ the full encoder-decoder transformer architecture [36] for conditioned crystal generation. The first step of property-guided generation could take into account of synthesizability and stability of materials. In the future, proposing materials with multiple properties is a promising direction for material generative models. To achieve that goal, curating a high-quality material dataset with either computed or measured properties is crucial. Needless to say, to close the loop of material discovery, one will need to carry out experimental verification for the generated materials.
Acknowledgements.
We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Qi Yang, Shigang Ou, Xinyang Dong, and Quansheng Wu for useful discussions. This project is supported by the National Natural Science Foundation of China under Grants No. T2225018, No. 92270107, No. 12188101, No. T2121001, and No. 12034009 and the Strategic Priority Research Program of Chinese Academy of Sciences under Grants No. XDB0500000 and No. XDB30000000.References
- Woodley and Catlow [2008] S. M. Woodley and R. Catlow, Crystal structure prediction from first principles, Nature Materials 7, 937 (2008).
- Oganov et al. [2019] A. R. Oganov, C. J. Pickard, Q. Zhu, and R. J. Needs, Structure prediction drives materials discovery, Nature Reviews Materials 4, 331 (2019).
- Gómez-Bombarelli et al. [2018] R. Gómez-Bombarelli, J. N. Wei, D. Duvenaud, J. M. Hernández-Lobato, B. Sánchez-Lengeling, D. Sheberla, J. Aguilera-Iparraguirre, T. D. Hirzel, R. P. Adams, and A. Aspuru-Guzik, Automatic chemical design using a data-driven continuous representation of molecules, ACS Central Science 4, 268 (2018), pMID: 29532027.
- Sanchez-Lengeling and Aspuru-Guzik [2018] B. Sanchez-Lengeling and A. Aspuru-Guzik, Inverse molecular design using machine learning: Generative models for matter engineering, Science 361, 360 (2018).
- Merchant et al. [2023] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80 (2023).
- Chen et al. [2024] C. Chen, D. T. Nguyen, S. J. Lee, N. A. Baker, A. S. Karakoti, L. Lauw, C. Owen, K. T. Mueller, B. A. Bilodeau, V. Murugesan, and M. Troyer, Accelerating computational materials discovery with artificial intelligence and cloud high-performance computing: from large-scale screening to experimental validation, (2024), arXiv:2401.04070 [cond-mat.mtrl-sci] .
- Achiam et al. [2023] J. Achiam, S. Adler, S. Agarwal, L. Ahmad, I. Akkaya, F. L. Aleman, D. Almeida, J. Altenschmidt, S. Altman, S. Anadkat, et al., Gpt-4 technical report, (2023), arXiv:2303.08774 [cs.CL] .
- Ramesh et al. [2021] A. Ramesh, M. Pavlov, G. Goh, S. Gray, C. Voss, A. Radford, M. Chen, and I. Sutskever, Zero-shot text-to-image generation, in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 139, edited by M. Meila and T. Zhang (PMLR, 2021) pp. 8821–8831.
- Rombach et al. [2022] R. Rombach, A. Blattmann, D. Lorenz, P. Esser, and B. Ommer, High-resolution image synthesis with latent diffusion models, in Proceedings of the IEEE/CVF conference on computer vision and pattern recognition (2022) pp. 10684–10695.
- Ingraham et al. [2023] J. B. Ingraham, M. Baranov, Z. Costello, K. W. Barber, W. Wang, A. Ismail, V. Frappier, D. M. Lord, C. Ng-Thow-Hing, E. R. Van Vlack, S. Tie, V. Xue, S. C. Cowles, A. Leung, J. V. Rodrigues, C. L. Morales-Perez, A. M. Ayoub, R. Green, K. Puentes, F. Oplinger, N. V. Panwar, F. Obermeyer, A. R. Root, A. L. Beam, F. J. Poelwijk, and G. Grigoryan, Illuminating protein space with a programmable generative model, Nature 623, 1070 (2023).
- Xie et al. [2021] T. Xie, X. Fu, O.-E. Ganea, R. Barzilay, and T. Jaakkola, Crystal diffusion variational autoencoder for periodic material generation, (2021), arXiv:2110.06197 [cs.LG] .
- Luo et al. [2023] Y. Luo, C. Liu, and S. Ji, Towards symmetry-aware generation of periodic materials, (2023), arXiv:2307.02707 [cs.LG] .
- Jiao et al. [2023] R. Jiao, W. Huang, P. Lin, J. Han, P. Chen, Y. Lu, and Y. Liu, Crystal structure prediction by joint equivariant diffusion, (2023), arXiv:2309.04475 [cond-mat.mtrl-sci] .
- Zheng et al. [2023] S. Zheng, J. He, C. Liu, Y. Shi, Z. Lu, W. Feng, F. Ju, J. Wang, J. Zhu, Y. Min, H. Zhang, S. Tang, H. Hao, P. Jin, C. Chen, F. Noé, H. Liu, and T.-Y. Liu, Towards predicting equilibrium distributions for molecular systems with deep learning, (2023), arXiv:2306.0544 [physics.chem-ph] .
- Yang et al. [2023] M. Yang, K. Cho, A. Merchant, P. Abbeel, D. Schuurmans, I. Mordatch, and E. D. Cubuk, Scalable diffusion for materials discovery, (2023), arXiv:2311.09235 [cs.LG] .
- Zeni et al. [2023] C. Zeni, R. Pinsler, D. Zügner, A. Fowler, M. Horton, X. Fu, S. Shysheya, J. Crabbé, L. Sun, J. Smith, et al., Mattergen: a generative model for inorganic materials design, (2023), arXiv:2312.03687 [cond-mat.mtrl-sci] .
- Xiao et al. [2023] H. Xiao, R. Li, X. Shi, Y. Chen, L. Zhu, X. Chen, and L. Wang, An invertible, invariant crystal representation for inverse design of solid-state materials using generative deep learning, Nature Communications 14, 7027 (2023).
- Flam-Shepherd and Aspuru-Guzik [2023] D. Flam-Shepherd and A. Aspuru-Guzik, Language models can generate molecules, materials, and protein binding sites directly in three dimensions as xyz, cif, and pdb files, (2023), arXiv:2305.05708 [cs.LG] .
- Antunes et al. [2023] L. M. Antunes, K. T. Butler, and R. Grau-Crespo, Crystal structure generation with autoregressive large language modeling, (2023), arXiv:2307.04340 [cond-mat.mtrl-sci] .
- Gruver et al. [2024] N. Gruver, A. Sriram, A. Madotto, A. G. Wilson, C. L. Zitnick, and Z. Ulissi, Fine-tuned language models generate stable inorganic materials as text, (2024), arXiv:2402.04379 [cs.LG] .
- Luo et al. [2024] X. Luo, Z. Wang, P. Gao, J. Lv, Y. Wang, C. Chen, and Y. Ma, Deep learning generative model for crystal structure prediction, (2024), arXiv:2403.10846 [cond-mat.mtrl-sci] .
- Ye et al. [2024] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, (2024), arXiv:2403.12478 [cond-mat.mtrl-sci] .
- Zhu et al. [2023] R. Zhu, W. Nong, S. Yamazaki, and K. Hippalgaonkar, Wycryst: Wyckoff inorganic crystal generator framework, (2023), arXiv:2311.17916 [cond-mat.mtrl-sci] .
- AI4Science et al. [2023] M. AI4Science, A. Hernandez-Garcia, A. Duval, A. Volokhova, Y. Bengio, D. Sharma, P. L. Carrier, M. Koziarski, and V. Schmidt, Crystal-gfn: sampling crystals with desirable properties and constraints, (2023), arXiv:2310.04925 [cs.LG] .
- Nguyen et al. [2024] T. M. Nguyen, S. A. Tawfik, T. Tran, S. Gupta, S. Rana, and S. Venkatesh, Hierarchical GFlownet for crystal structure generation (2024).
- Jiao et al. [2024] R. Jiao, W. Huang, Y. Liu, D. Zhao, and Y. Liu, Space group constrained crystal generation, (2024), arXiv:2402.03992 [cs.LG] .
- Oganov and Glass [2006] A. R. Oganov and C. W. Glass, Crystal structure prediction using ab initio evolutionary techniques: Principles and applications, The Journal of chemical physics 124 (2006).
- Davies et al. [2016] D. W. Davies, K. T. Butler, A. J. Jackson, A. Morris, J. M. Frost, J. M. Skelton, and A. Walsh, Computational screening of all stoichiometric inorganic materials, Chem 1, 617 (2016).
- Glazer et al. [2012] M. Glazer, G. Burns, and A. Glazer, Space Groups for Solid State Scientists (Elsevier Science, 2012).
- Urusov and Nadezhina [2009] V. S. Urusov and T. N. Nadezhina, Frequency distribution and selection of space groups in inorganic crystal chemistry, Journal of Structural Chemistry 50, 22 (2009).
- Marsh [1999] R. E. Marsh, P1 or P1¯? Or something else?, Acta Crystallographica Section B 55, 931 (1999).
- [32] See https://github.com/deepmodeling/CrystalFormer for code and model checkpoint.
- Zhao et al. [2020] Y. Zhao, Y. Cui, Z. Xiong, J. Jin, Z. Liu, R. Dong, and J. Hu, Machine learning-based prediction of crystal systems and space groups from inorganic materials compositions, ACS Omega 5, 3596 (2020), pMID: 32118175.
- Wang et al. [2023a] D.-Y. Wang, H.-F. Lv, and X.-J. Wu, Crystallographic groups prediction from chemical composition via deep learning, Chinese Journal of Chemical Physics 36, 66 (2023a).
- Hahn et al. [1983] T. Hahn, U. Shmueli, and J. W. Arthur, International tables for crystallography, Vol. 1 (Reidel Dordrecht, 1983).
- Vaswani et al. [2017] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. u. Kaiser, and I. Polosukhin, Attention is all you need, in Advances in Neural Information Processing Systems, Vol. 30, edited by I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Curran Associates, Inc., 2017).
- Wirnsberger et al. [2020] P. Wirnsberger, A. J. Ballard, G. Papamakarios, S. Abercrombie, S. Racanière, A. Pritzel, D. J. Rezende, and C. Blundell, Targeted free energy estimation via learned mappings, Journal of Chemical Physics 153, 144112 (2020).
- Note [1] We follow the convention that capital letters appear after lower case letters. This handles the edge case of the No. 47 space group whose Wyckoff positions reach ’A’. This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’.
- [39] OpenAI, Using logit bias to alter token probability with the openai api, OpenAI Help Center, retrieved March 14, 2024.
- Xie et al. [2023] H. Xie, L. Zhang, and L. Wang, of two-dimensional electron gas: A neural canonical transformation study, SciPost Phys. 14, 154 (2023).
- Parthé and Gelato [1984] E. Parthé and L. M. Gelato, The standardization of inorganic crystal-structure data, Acta Crystallographica Section A 40, 169 (1984).
- Glawe et al. [2016] H. Glawe, A. Sanna, E. K. U. Gross, and M. A. L. Marques, The optimal one dimensional periodic table: a modified pettifor chemical scale from data mining, New Journal of Physics 18, 093011 (2016).
- Zhou et al. [2018] Q. Zhou, P. Tang, S. Liu, J. Pan, Q. Yan, and S.-C. Zhang, Learning atoms for materials discovery, Proceedings of the National Academy of Sciences 115, E6411 (2018).
- Chen et al. [2019] C. Chen, W. Ye, Y. Zuo, C. Zheng, and S. P. Ong, Graph networks as a universal machine learning framework for molecules and crystals, Chemistry of Materials 31, 3564 (2019).
- Zhang et al. [2022] D. Zhang, H. Bi, F.-Z. Dai, W. Jiang, L. Zhang, and H. Wang, Dpa-1: Pretraining of attention-based deep potential model for molecular simulation, (2022), arXiv:2208.08236 [physics.chem-ph] .
- Wang et al. [2022] A. Y.-T. Wang, M. S. Mahmoud, M. Czasny, and A. Gurlo, Crabnet for explainable deep learning in materials science: bridging the gap between academia and industry, Integrating Materials and Manufacturing Innovation 11, 41 (2022).
- Wang et al. [2021a] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021a).
- Pauling [1929] L. Pauling, The principles determining the structure of complex ionic crystals, Journal of the American Chemical Society 51, 1010 (1929).
- Goldschmidt [1929] V. Goldschmidt, Crystal structure and chemical constitution, Transactions of the Faraday Society 25, 253 (1929).
- Pettifor [1988] D. Pettifor, Structure maps for. pseudobinary and ternary phases, Materials Science and Technology 4, 675 (1988).
- Fischer et al. [2006] C. C. Fischer, K. J. Tibbetts, D. Morgan, and G. Ceder, Predicting crystal structure by merging data mining with quantum mechanics, Nature Materials 5, 641 (2006).
- Allahyari and Oganov [2020] Z. Allahyari and A. R. Oganov, Coevolutionary search for optimal materials in the space of all possible compounds, npj Computational Materials 6, 55 (2020).
- Court et al. [2020] C. J. Court, B. Yildirim, A. Jain, and J. M. Cole, 3-d inorganic crystal structure generation and property prediction via representation learning, Journal of Chemical Information and Modeling 60, 4518 (2020), pMID: 32866381.
- Davies et al. [2019] D. W. Davies, K. T. Butler, A. J. Jackson, J. M. Skelton, K. Morita, and A. Walsh, Smact: Semiconducting materials by analogy and chemical theory, Journal of Open Source Software 4, 1361 (2019).
- Gebauer et al. [2019] N. Gebauer, M. Gastegger, and K. Schütt, Symmetry-adapted generation of 3d point sets for the targeted discovery of molecules, in Advances in Neural Information Processing Systems, Vol. 32 (Curran Associates, Inc., 2019).
- [56] https://github.com/sparks-baird/matbench-genmetrics.
- Goodfellow et al. [2016] I. Goodfellow, Y. Bengio, and A. Courville, Deep Learning (MIT Press, 2016) http://www.deeplearningbook.org.
- Sun et al. [2016] W. Sun, S. T. Dacek, S. P. Ong, G. Hautier, A. Jain, W. D. Richards, A. C. Gamst, K. A. Persson, and G. Ceder, The thermodynamic scale of inorganic crystalline metastability, Science Advances 2, e1600225 (2016).
- Note [2] Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No. 221).
- Ong et al. [2013] S. P. Ong, W. D. Richards, A. Jain, G. Hautier, M. Kocher, S. Cholia, D. Gunter, V. L. Chevrier, K. A. Persson, and G. Ceder, Python materials genomics (pymatgen): A robust, open-source python library for materials analysis, Computational Materials Science 68, 314 (2013).
- Jain et al. [2013] A. Jain, S. P. Ong, G. Hautier, W. Chen, W. D. Richards, S. Dacek, S. Cholia, D. Gunter, D. Skinner, G. Ceder, and K. A. Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
- Pickard and Needs [2011] C. J. Pickard and R. J. Needs, Ab initio random structure searching, Journal of Physics: Condensed Matter 23, 053201 (2011).
- Wang et al. [2012] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Calypso: A method for crystal structure prediction, Computer Physics Communications 183, 2063 (2012).
- Lyakhov et al. [2013] A. O. Lyakhov, A. R. Oganov, H. T. Stokes, and Q. Zhu, New developments in evolutionary structure prediction algorithm uspex, Computer Physics Communications 184, 1172 (2013).
- Avery and Zurek [2017] P. Avery and E. Zurek, Randspg: An open-source program for generating atomistic crystal structures with specific spacegroups, Computer Physics Communications 213, 208 (2017).
- Fredericks et al. [2021] S. Fredericks, K. Parrish, D. Sayre, and Q. Zhu, Pyxtal: A python library for crystal structure generation and symmetry analysis, Computer Physics Communications 261, 107810 (2021).
- Cheng et al. [2022] G. Cheng, X.-G. Gong, and W.-J. Yin, Crystal structure prediction by combining graph network and optimization algorithm, Nature Communications 13, 1492 (2022).
- Wang et al. [2023b] H.-C. Wang, J. Schmidt, M. A. L. Marques, L. Wirtz, and A. H. Romero, Symmetry-based computational search for novel binary and ternary 2d materials, 2D Materials 10, 035007 (2023b).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
- Zhang et al. [2023] D. Zhang, X. Liu, X. Zhang, C. Zhang, C. Cai, H. Bi, Y. Du, X. Qin, J. Huang, B. Li, Y. Shan, J. Zeng, Y. Zhang, S. Liu, Y. Li, J. Chang, X. Wang, S. Zhou, J. Liu, X. Luo, Z. Wang, W. Jiang, J. Wu, Y. Yang, J. Yang, M. Yang, F.-Q. Gong, L. Zhang, M. Shi, F.-Z. Dai, D. M. York, S. Liu, T. Zhu, Z. Zhong, J. Lv, J. Cheng, W. Jia, M. Chen, G. Ke, W. E, L. Zhang, and H. Wang, Dpa-2: Towards a universal large atomic model for molecular and material simulation, (2023), arXiv:2312.15492 [physics.chem-ph] .
- Batatia et al. [2024] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, M. Avaylon, W. J. Baldwin, F. Berger, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, F. Falcioni, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, P. Grigorev, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, J. R. Kermode, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, B. X. Shi, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, T. D. Swinburne, J. Tilly, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills, and G. Csányi, A foundation model for atomistic materials chemistry, (2024), arXiv:2401.00096 [physics.chem-ph] .
- Ren et al. [2022] Z. Ren, S. I. P. Tian, J. Noh, F. Oviedo, G. Xing, J. Li, Q. Liang, R. Zhu, A. G. Aberle, S. Sun, et al., An invertible crystallographic representation for general inverse design of inorganic crystals with targeted properties, Matter 5, 314 (2022).
- Zhao et al. [2021] Y. Zhao, M. Al-Fahdi, M. Hu, E. M. Siriwardane, Y. Song, A. Nasiri, and J. Hu, High-throughput discovery of novel cubic crystal materials using deep generative neural networks, Advanced Science 8, 2100566 (2021).
- Ahmad and Cai [2022] R. Ahmad and W. Cai, Free energy calculation of crystalline solids using normalizing flows, Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
- Wirnsberger et al. [2022] P. Wirnsberger, G. Papamakarios, B. Ibarz, S. Racanière, A. J. Ballard, A. Pritzel, and C. Blundell, Normalizing flows for atomic solids, Machine Learning: Science and Technology 3, 025009 (2022).
- Köhler et al. [2023] J. Köhler, M. Invernizzi, P. De Haan, and F. Noé, Rigid body flows for sampling molecular crystal structures, (2023), arXiv:2301.11355 [cs.LG] .
- Gebauer et al. [2018] N. W. Gebauer, M. Gastegger, and K. T. Schütt, Generating equilibrium molecules with deep neural networks, (2018), arXiv:1810.11347 [stat.ML] .
- Gebauer et al. [2022] N. W. A. Gebauer, M. Gastegger, S. S. P. Hessmann, K.-R. Müller, and K. T. Schütt, Inverse design of 3d molecular structures with conditional generative neural networks, Nature Communications 13, 973 (2022).
- Yao et al. [2023] S. Yao, D. Yu, J. Zhao, I. Shafran, T. L. Griffiths, Y. Cao, and K. Narasimhan, Tree of thoughts: Deliberate problem solving with large language models, (2023), arXiv:2305.10601 [cs.CL] .
- Jain and Bligaard [2018] A. Jain and T. Bligaard, Atomic-position independent descriptor for machine learning of material properties, Phys. Rev. B 98, 214112 (2018).
- Goodall et al. [2022] R. E. A. Goodall, A. S. Parackal, F. A. Faber, R. Armiento, and A. A. Lee, Rapid discovery of stable materials by coordinate-free coarse graining, Science Advances 8, eabn4117 (2022).
- Yan et al. [2022] K. Yan, Y. Liu, Y. Lin, and S. Ji, Periodic graph transformers for crystal material property prediction, in Advances in Neural Information Processing Systems, Vol. 35, edited by S. Koyejo, S. Mohamed, A. Agarwal, D. Belgrave, K. Cho, and A. Oh (Curran Associates, Inc., 2022) pp. 15066–15080.
- Taniai et al. [2024] T. Taniai, R. Igarashi, Y. Suzuki, N. Chiba, K. Saito, Y. Ushiku, and K. Ono, Crystalformer: Infinitely connected attention for periodic structure encoding, (2024), arXiv:2403.11686 [cs.LG] .
- Rubungo et al. [2023] A. N. Rubungo, C. Arnold, B. P. Rand, and A. B. Dieng, Llm-prop: Predicting physical and electronic properties of crystalline solids from their text descriptions, (2023), arXiv:2310.14029 [cs.CL] .
- Maddox [1988] J. Maddox, Crystals from first principles, Nature 335, 201 (1988).
- Brown et al. [2020] T. Brown, B. Mann, N. Ryder, M. Subbiah, J. D. Kaplan, P. Dhariwal, A. Neelakantan, P. Shyam, G. Sastry, A. Askell, et al., Language models are few-shot learners, Advances in neural information processing systems 33, 1877 (2020).
- Kaplan et al. [2020] J. Kaplan, S. McCandlish, T. Henighan, T. B. Brown, B. Chess, R. Child, S. Gray, A. Radford, J. Wu, and D. Amodei, Scaling laws for neural language models, (2020), arXiv:2001.08361 [cs.LG] .
- Frey et al. [2023] N. C. Frey, R. Soklaski, S. Axelrod, S. Samsi, R. Gómez-Bombarelli, C. W. Coley, and V. Gadepally, Neural scaling of deep chemical models, Nature Machine Intelligence 5, 1297 (2023).
- Stokes and Hatch [1989] H. T. Stokes and D. M. Hatch, Isotropy Subgroups of the 230 Crystallographic Space Groups (WORLD SCIENTIFIC, 1989).
- Müller [2013] U. Müller, Symmetry Relationships between Crystal Structures: Applications of Crystallographic Group Theory in Crystal Chemistry, International Union of Crystallography Texts on Crystallography (OUP Oxford, 2013).
- Cox and White [2022] S. Cox and A. D. White, Symmetric molecular dynamics, Journal of Chemical Theory and Computation 18, 4077 (2022).
- Lai et al. [2024] Q. Lai, L. Yao, Z. Gao, S. Liu, H. Wang, S. Lu, D. He, L. Wang, C. Wang, and G. Ke, End-to-end crystal structure prediction from powder x-ray diffraction, (2024), arXiv:2401.03862 [physics.chem-ph] .
- Wang et al. [2021b] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021b).
- Schmidt et al. [2022a] J. Schmidt, N. Hoffmann, H.-C. Wang, P. Borlido, P. J. M. A. Carriço, T. F. T. Cerqueira, S. Botti, and M. A. L. Marques, Large-scale machine-learning-assisted exploration of the whole materials space (2022a), arXiv:2210.00579 [cond-mat.mtrl-sci] .
- Schmidt et al. [2022b] J. Schmidt, H.-C. Wang, T. F. T. Cerqueira, S. Botti, and M. A. L. Marques, A dataset of 175k stable and metastable materials calculated with the pbesol and scan functionals, Scientific Data 9, 64 (2022b).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized Gradient Approximation Made Simple, Phys. Rev. Lett. 77, 3865 (1996).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
Appendix A More details of CrystalFormer
To recap, the space group information plays a key role in the architecture, training, and sampling of CrystalFormer. First of all, the vector embedding of space group number controls all subsequent outputs in the transformer that determines the Wyckoff letters, chemical species, fractional coordinates, and lattice parameters. Second, the information of the space group and Wyckoff letter are used to select active components in the fractional coordinates and lattice parameters in the loss function during training. Lastly, the space group determines the concrete meaning of in terms of multiplicities and fractional coordinates. This information is used to place right amount of atoms to appropriate locations during sampling.
A.1 Model architectures
Algorithm 1 summarized the model architecture of CrystalFormer. Training the model for 3,800 epochs with the hyperparameters shown in Table S1 takes about 13 hours on a single A100 GPU.
| Hyperparameters | Value | Remarks |
| The length of atom sequence including the padding atoms | 21 | |
| Number of chemical species | 119 | ’H’ to ’Og’, plus padding atom |
| Number of possible Wyckoff letters | 28 | ’a-z’+’A’, plus padding atom |
| Number of modes in von-Mises mixture distribution | 16 | |
| Number of modes in lattice Gaussian mixture distribution | 16 | |
| Hidden layer dimension for the composite type of the first atom | 256 | |
| Transformer number of layers | 16 | |
| Transformer number of heads | 16 | |
| Transformer key size | 64 | |
| Transformer model size | 32 | |
| Embedding dimension of discrete input | 32 | |
| Number of Fourier frequency | 5 | |
| Learning rate | 0.0001 | |
| Learning rate decay | 0.0 | |
| Weight decay | 0.0 | |
| Clip grad | 1.0 | |
| Batch Size | 100 | |
| Optimizer | Adam | |
| Dropout rate | 0.5 | |
| Total number of parameters: 4840295 |
A.2 Sampling algorithm
Algorithm 2 summarizes the sampling method of CrystalFormer. It takes 520 seconds to generate a batch size 13,000 crystal samples on a single A100 GPU, which translates to a generation speed 40 millisecond per sample.
Appendix B Stable crystal samples
Table S2 list generated samples in the space group (No. 225) with . Three of them (InNd, InTb, and EuTl) actually fall into the space group (No. 221) due to merging of ”4a” and ”4b” Wyckoff positions that are occupied by the same elements. Overall, out of 1000 samples generated by CrystalFormer, we found 10 stable materials that are not contained in the Materials Project database. Extending the search to more recent material databases, we found match of AcCuRh, AcHgIn, AcAgSi and AuInPm in the WBM dataset [93], AcHgIn, AcAgSi, InPmTl, PdPmZn, IrLuPm and AuInPm in the Alexandria dataset [94, 95]. Neverthless, we did not find AcBiK, InTb, and CaPaRu to be present in any of these dataset.
To estimate the energy above hull, the DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [96] and all-electron projector-augmented wave method [97], as implemented in the VASP code [98]. All parameters of the calculations including settings of PBE functional, Hubbard U corrections, and ferromagnetic initialization are chosen to be consistent with Materials Project by using of MPRelaxSet function in pymatgen [60] A double relaxation strategy was employed. The maximum optimization ionic step and the maximum running time were constrained to 150 steps and 20 hours, respectively. All structures containing Yb element are ignored when calculating energy above hull due to they are unavailable from the Materials Project at the time of writing.
| In MP | Not in MP | ||||
| Formula | (eV/atom) | Lattice constant () | Formula | (eV/atom) | Lattice constant () |
| PdPtTm | -0.00164 | 6.90 (6.86) | AcCuRh | -0.14248 | 7.47 |
| LiPtZr | -0.01668 | 6.43 (6.39) | AcAgSi | -0.04364 | 7.81 |
| CdLiPb | -0.00313 | 6.82 (6.79) | IrLuPm | -0.0289 | 6.89 |
| LiNdPb | -0.05703 | 7.05 (7.03) | InPmTl | -0.01318 | 7.73 |
| BeOsSi | -0.00779 | 5.77 (5.74) | PdPmZn | -0.0843 | 7.27 |
| CeAs | -0.0073 | 5.99 (6.09) | AuInPm | -0.0568 | 7.70 |
| Xe | -0.00067 | 7.45 (6.66) | AcBiK | -0.20215 | 8.37 |
| InLiY | -0.17504 | 7.54 | InTb | -0.26278 | 7.95 |
| InNd | -0.21393 | 8.35 | AcHgIn | -0.40448 | 8.66 |
| BeC | -0.01227 | 4.33 (4.32) | CaPaRu | -0.09344 | 6.80 |
| EuTl | -0.1054 | 7.90 | |||

### v2

@@
 We introduce CrystalFormer, a transformer-based autoregressive model specifically designed for space group-controlled generation of crystalline materials.
-The space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials.
+The incorporation of space group symmetry significantly simplifies the crystal space, which is crucial for data and compute efficient generative modeling of crystalline materials.
 Leveraging the prominent discrete and sequential nature of the Wyckoff positions, CrystalFormer learns to generate crystals by directly predicting the species and locations of symmetry-inequivalent atoms in the unit cell.
-Our results demonstrate that CrystalFormer matches state-of-the-art performance on standard benchmarks for both validity, novelty, and stability of the generated crystalline materials.
-Our analysis also shows that CrystalFormer ingests sensible solid-state chemistry information from data for generative modeling.
-The CrystalFormer unifies symmetry-based structure search and generative pre-training in the realm of crystalline materials.
+We demonstrate the advantages of CrystalFormer in standard tasks such as symmetric structure initialization and element substitution compared to conventional methods implemented in popular crystal structure prediction software.
+Moreover, we showcase the application of CrystalFormer of property-guided materials design in a plug-and-play manner.
+Our analysis shows that CrystalFormer ingests sensible solid-state chemistry knowledge and heuristics by compressing the material dataset, thus enabling systematic exploration of crystalline materials.
 The simplicity, generality, and flexibility of CrystalFormer position it as a promising architecture to be the foundational model of the entire crystalline materials space, heralding a new era in materials modeling and discovery.
+Contents
+- I Introduction
+- II CrystalFormer
+- III CrystalFormer learns chemical intuition by compressing materials database
+- IV Applications
+- V Related works
+- VI Outlook
+- A More details of CrystalFormer
+- B Validity and novelty of generated samples
+- C Discovered crystal samples with symmetric structure initialization
+- D Discovered crystal samples with element substitution
+- E Details of plug-and-play materials design
 I Introduction
@@
 Generative machine learning, in particular, has been a promising step for matter inverse design [3, 4] which goes beyond machine learning accelerated structure search [5] and property screening [6].
-Generative models learn the underlying distribution of training data and generated new samples from the learned distribution.
-In addition, the generation process can also be controlled by prompts such as desired material properties or experiment observations.
+Generative models learn the underlying distribution of training data and generate new samples from the learned distribution.
+In addition, the generation process can also be controlled by conditions such as desired material properties or experiment observations.
 Amazing programming abilities of generative models have been demonstrated in large language model [7], text-to-image generation [8, 9], and protein design [10].
 It is anticipated that generative model-based approaches will introduce groundbreaking changes to the traditional workflows of material discovery.
-A generative pretrained foundation model for crystalline materials is a key step towards such lofty goal.
-However, despite of intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains.
-Simply scaling the compute and model size of the current crystal generative model may not be enough because the amount of high-quality data for crystalline materials is much less than compared to language and image domains.
-Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [12, 23, 24, 25, 26].
+A generative pre-trained foundation model for crystalline materials is a key step towards such a lofty goal.
+However, despite intensive efforts [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the current generative models for crystalline materials fall short to match the success of other domains.
+Simply scaling the compute and model size of the current crystal generative model may not be feasible because the amount of high-quality data for crystalline materials is much less than compared to language and image domains.
+Therefore, leveraging the inherent inductive biases specific to crystalline structures for more data-efficient generative modeling is essential, as has been pursued in some of recent works [23, 24, 25, 26].
 | world | With space group symmetry |
-The space group symmetry is arguably the most important inductive bias in the modeling of crystalline materials, which is the joint outcome of the rotational and translational symmetry in space.
+The space group symmetry due to the joint outcome of the rotational and translational symmetry in space is arguably the most important inductive bias in the modeling of crystalline materials.
 There are in total 230 space groups [29] for three-dimensional crystal structures.
 Nature exhibits a preference for symmetric crystal structures, a tendency that may be attributed to the symmetry inherent in the interatomic interactions, which, in turn, are governed by the fundamental forces acting between elementary particles.
-As a result, the appearance of crystalline materials in the the first and the least symmetric space group is rare [30], with many instances potentially being misclassified.
-[31].
+As a result, the appearance of crystalline materials in the first and the least symmetric space group is rare [30], with many instances potentially even being misclassified [31].
+Failing to match the space group distribution of nature in machine learning-generated materials is regarded as a matter of serious concern [32].
 Space group symmetry imposes significant constraints on a crystal.
-At its core, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the dimensions and angles of the crystal’s unit cell.
-Moreover, the symmetry operations associated with a given space group ensure that identical atoms are consistently mapped onto each other across the crystal, preserving its structural integrity.
-This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the total number of atoms in each unit cell.
-A key concept in understanding these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group.
+First of all, the space group identifies the crystal system to which a crystal belongs, thereby limiting the permissible values for the lattice parameters that define the length and angles of the crystal’s unit cell.
+Moreover, the symmetry operations associated with a given space group ensure that symmetry equivalent atoms are consistently mapped among themselves in the crystal.
+This requirement enforces strict conditions regarding the types of chemical elements present, their specific locations within the crystal, and the number of each chemical species in the unit cell.
+A key concept to express these constraints is the Wyckoff positions, which delineate unique areas within a unit cell that are defined by the symmetry operations of the crystal’s space group.
 These positions are represented as fractional coordinates, enabling precise definition relative to the unit cell’s axes.
@@
 For example, the top row of the table in Fig.
-1(a) shows a general position that can be mapped to 36 positions under the symmetry operations of the space group.
+1(a) contains the general position that can be mapped to 36 positions under the symmetry operations of the space group.
 Nature tends to place atoms in those special Wyckoff positions at the bottom of the table.
-For example, we highlight the occupied Wyckoff positions of calcite (CaCO) crystal in Fig.
+For example, we highlight the occupied Wyckoff positions of calcite (CaCO3) crystal in Fig.
 1, associated with the space group.
-One sees that the Wyckoff letter ’6a’ and ’6b’ deterministicly define the locations of the carbon and calcium atoms within the unit cell.
+One sees that the Wyckoff letter ’6a’ and ’6b’ deterministically define the locations of the carbon and calcium atoms within the unit cell.
 In addition, it follows that , and as the space group belongs to the trigonal crystal system.
-Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO structure: the coordinate of oxygen atom and the lattice constants and .
+Ultimately, despite having 30 atoms in the unit cell, there are only three continuous degrees of freedom for the CaCO3 structure: the x-coordinate of oxygen atom and the lattice constants and .
 All other information about the crystal structure can be specified via discrete data such as the Wyckoff letters and chemical species.
 The prominent discrete and sequential features illustrated in Figure 1 are ubiquitous in crystalline materials.
-The Wyckoff positions not only specify possible locations of atoms in the unit cell, their associated multiplicities also put strong constraints on the number of atoms.
+The Wyckoff positions not only specify possible locations of atoms in the unit cell, but their associated multiplicities also put strong constraints on the number of atoms.
 Therefore, space group symmetry significantly reduces the degrees of freedom of crystalline materials.
@@
 For example, the performance of the generative model quickly deteriorates as the number of atoms increases due to it is challenging to generate highly symmetric crystal structures [16].
+On the other hand, statistical analysis shows that the Wyckoff sequences of known inorganic compounds [33] are far from being exhausted, implying there are statistical correlations to be exploited to compress the materials database.
 In this paper, we introduce CrystalFormer, an autoregressive transformer for generative modeling of crystalline materials.
 CrystalFormer models the joint probability distribution of Wyckoff positions, chemical species, and lattice parameters of crystals with a given space group.
-By treating the Wyckoff positions as the first class citizen in the model, CrystalFormer seamlessly integrates the space group symmetry into crystal probabilistic modeling.
-The space group-informed transformer exploits the fundamental inductive bias of crystals to greatly constrain and simplify the generation process.
+By treating the Wyckoff positions as the first class citizen in the model, CrystalFormer seamlessly and rigorously integrates the space group symmetry into crystal probabilistic modeling.
 As shown in Table 1, explicit modeling of the Wyckoff positions greatly reduces the space of crystalline materials.
-We benchmark the CrystalFormer model on standard crystal generation tasks and show it exhibits great efficiency and generalization ability compared to existing models.
-Finally, we discuss its potential applications in novel materials discovery.
+The space group-informed transformer exploits this fundamental inductive bias to greatly simplify the learning and generation of crystals.
+The organization of the present paper is as follows.
+Section II introduces the model architecture, training, and sampling of the CrystalFormer.
+Section III reveals the chemical intuition encoded in the trained model by inspecting its weights and generated crystal samples.
+Such inspection also builds up an understanding of the strength of the model.
+Section IV demonstrates applications CrystalFormer for symmetric structure initialization and element substitution, where it exhibits great efficiency and generality compared to existing approaches.
+Furthermore, we showcase CrystalFormer’s ability for property-guided materials design in a plug-and-play manner.
+Section V puts the key contribution of present work in the broad context of crystal generative models.
+Finally, Section VI outlooks for possible extension and future of the present line of research.
 We have released the codes and trained model at [34].
 II CrystalFormer
+We will introduce the design, training, and sampling of the CrystalFormer model.
+II.1 Model
 To exploit the space group symmetry of the crystal, we focus on the Wyckoff positions of symmetry-inequivalent atoms.
-Wyckoff letters follows the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group.
+Wyckoff letters follow the alphabetical order, where ”a” stands for the positions with the highest order of symmetry for the given space group.
 Later letters in the alphabet indicate more general positions with reduced site symmetries.
 Note that the information of the space group number and Wyckoff letter fully determine the multiplicities.
-In cases where the atom positions are fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g.
-the -coordinate of the oxygen atoms in the CaCO example shown in Fig.
-1, To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially.
+In cases where the atom positions are not fully fixed by the Wyckoff letter, we will also consider the remaining fractional coordinates, e.g.
+the -coordinate of the oxygen atoms in the CaCO3 example shown in Fig.
+1.
+To generate crystals, one samples the Wyckoff letter, chemical element, and fractional coordinates of each atom sequentially.
 The sampling procedure starts from special higher symmetry sites with smaller multiplicities and then goes on to general lower symmetry regions with larger multiplicities.
@@
 For example, as shown in Fig.
-1(b) one has for CaCO.
+1(b) one has for CaCO3.
 Explicitly including the Wyckoff letter in the generative modeling is the key of the present work.
 Next, are the fractional coordinates of symmetrically inequivalent atoms.
-Lastly, denotes the lattice parameters of the conventional unit cell of the material.
+Lastly, denotes the lattice parameters of the conventional unit cell of the crystal.
 The central quantity to focus on is the conditional probability of a crystal given the space group number : .
 Since the space group is a fundamental characterization for crystalline materials, is a key control variable that greatly simplifies the distribution over the entire crystal materials space.
-In practical applications, the space group can be considered separately or even predicted [33, 34].
+In practical applications of crystal structure prediction and material design, the space group can either be considered separately as a control variable or predicted based on material composition [35, 36, 37, 38].
 We express the space group conditioned probability distribution of crystals as an autoregressive product of conditional probabilities
-At first sight, it may appear unnatural to employ an autoregressive model for crystals since there is no obvious order for atoms in the unit cell.
+At first sight, it may appear unnatural to employ an autoregressive model for crystals since there seems to be no obvious order for atoms in the unit cell.
 However, the sequential nature of Wyckoff positions suggests a natural way to arrange symmetrically inequivalent atoms in an alphabetical order of the Wyckoff letters.
-During sampling, the autoregressive model places the atoms one by one into the unit cell, starting from the special position with high symmetry to the general position with the lowest symmetry, see Fig.
-1(c).
+Following this key observation, we represent crystal data as sequences of space groups, Wyckoff letters, chemical species, and fractional coordinates of each symmetrically inequivalent atom.
+Together with the information lattice parameters, such sequence fully characterizes the compositional and structural information of crystalline material.
+Since statistical analysis reveals that anions are in less symmetric positions than cations for inorganic crystals [30], one would expect that anion atoms will typically appear after cation atoms in such a sequence.
+For example, CaCO3 is represented as a string ”167-a-C-0-0-1/4-b-Ca-0-0-0-e-O-0.257-0-1/4”.
+Autoregressive sampling of such a string means the model generates the crystal by placing the atoms sequentially into the unit cell, starting from the special position with high site symmetry to the general position with the lowest site symmetry, see Fig.
+1(d).
 We model the conditional probability of the Wyckoff letters and chemical species as categorical distributions.
-A mixture of von-Mises distribution is employed to model the conditional probability of the factional coordinates .
+On the other hand, we model the conditional probability of the factional coordinates as a mixture of von Mises distribution for continuous periodic variables.
 For Wyckoff positions with multiplicities greater than one, we only consider the first of fractional coordinates that appear in the international tables for crystallography [39].
 Lastly, we model the conditioned distribution of lattice parameters as a Gaussian mixture model.
-We build CrystalFormer using the vanilla decoder-only transformer [36] to model the probability distribution of crystalline materials Eq.
+We build CrystalFormer, an autoregressive transformer [40] to model the space group conditioned-probability distribution of crystalline materials Eq.
 (1).
-The transformer-based autoregressive model is a favorable choice to capture these prominent discrete characteristics of crystals.
-We feed atomistic features into an autoregressive transformer.
-The space group number is the first input to the transformer.
-The remaining inputs alternate between the Wyckoff letter, chemical species, and fractional coordinates of each atom.
+The space group number is the first input to CrystalFormer.
+The remaining inputs are the Wyckoff letter, chemical species, and fractional coordinates of each atom.
 One can go through the table of Fig.
@@
 We feed vector embeddings of the space group number, Wyckoff letter, and the chemical species input to the CrystalFormer.
-In particular, we concatenate the vector embedding of to all other inputs since it is the key control variable for the crystal generation.
-In addition, we have also provided the multiplicity of each Wyckoff position as an additional feature.
+In particular, we also concatenate the vector embedding of to all other inputs since it is the key control variable for the crystal generation.
+Moreover, we have also provided the multiplicity of each Wyckoff position as an additional feature.
 The multiplicity can be easily inferred from the space group and the Wyckoff letters.
-We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [37, 13] We pad the atom sequence up to a maximum length and take the shifted output as parameters of the conditional probability distribution Eq.
+We feed the fractional coordinates as Fourier features into the transformer so that the model preserves the periodicity of the unit cell [41, 13].
+We pad the atom sequence up to a maximum length and treat the output as parameters of the conditional probability distribution Eq.
 (1), see Fig.
@@
 At the location of the first padding atom, we predict the lattice parameters.
-We implement a number of constraints in the model to further reduce the phase space of the model.
+We implement a number of constraints in the model to further reduce its phase space.
 First, the Wyckoff letters should be valid for the given space group.
-For example, for the space group number 167 the Wyckoff letters go from ’a’ up to ’f’.
+For example, for the space group (No.
+167) the Wyckoff letters go from ’a’ up to ’f’.
 Second, we require that the Wyckoff letters follow alphabetical order in the sequence 111We follow the convention that capital letters appear after lower case letters.
-This handles the edge case of the No.
-47 space group whose Wyckoff positions reach ’A’.
-This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’..
-Lastly, the Wyckoff positions with no continuous degree of freedom (such as ’a’, ’b’, and ’d’ positions in the space group 167) can only be occupied once.
-Those constraints are implemented by tuning the logit biases of Wyckoff letters [39, 40].
-The design of CrystalFormer focuses mostly on the space group symmetries since we believe they are the most crucial ones for crystalline materials.
+This handles the edge case of the space group (No.
+47) whose Wyckoff letters used up 26 lowercase letters and reached ’A’ for the generic position.
+In addition, we use the letter ’X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g.
+of Fig.
+1(b)..
+Lastly, the Wyckoff positions with no free fractional coordinates (such as ’a’, ’b’, and ’d’ positions in the space group) can only be occupied once.
+Those constraints are implemented by setting the logit biases of Wyckoff letters to mask out invalid sequences [43, 44].
+The design of CrystalFormer focuses mostly on the space group symmetries which we believe to be the most important inductive bias for crystalline materials.
 This design decision significantly impacts the treatment of other symmetries.
@@
 Note that in a crystal environment, the same type of atoms occupying different Wyckoff positions could be regarded as distinguished particles as they generally have different site symmetry.
-Lastly, the periodicity of the fractional coordinates is respected since they are treated as variables under the von-Mises distribution.
+Lastly, the periodicity of the fractional coordinates is respected in CrystalFormer since they are treated as periodic variables under the von Mises distribution.
+II.2 Training
 The CrystalFormer is trained by minimizing the negative log-likelihood over training dataset
@@
 , ) do not contribute to the loss function.
-To sample crystals from the CrystalFormer, one needs to specify a space group number and a number of possible chemical elements.
+In the present work, we train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell.
+The training dataset contains 27136 crystal structures.
+The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category.
+On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalanced distribution of crystals over space groups in nature [30].
+In fact, there is no training data in 61 out of 230 space groups as shown in Fig.
+S2.
+Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature, see appendix B.
+Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters.
+We select the model checkpoint with the lowest total validation loss to generate crystal samples.
+II.3 Sampling
+To sample crystals from the CrystalFormer, one needs to specify a space group number and a list of possible chemical elements.
 The CrystalFormer samples the atoms one by one, starting from more symmetric specific positions with lower multiplicities till less symmetric general positions with larger multiplicities.
 We use the information of the space group and Wyckoff letter to control the sampling of fractional coordinates.
-By applying the symmetry projection to the sampled fractional coordinate, one ensures the generated fractional coordinates are compatible with the Wyckoff positions.
+By applying the symmetry projection to the sampled fractional coordinate, one rectifies it and ensures the generated fractional coordinates are compatible with the Wyckoff positions.
 One can also mask out the logits of chemical species so that only a number of selected elements will be sampled.
-The number of symmetrically inequalivent atoms may fluctuate in the sampling procedure.
+The number of symmetrically inequivalent atoms may fluctuate in the sampling procedure.
 Once one has sampled a padding atom, the model predicts the lattice parameters under the space group constraint.
@@
 With we will draw samples from a sharper distribution, while gives more diversity in the generated samples.
-III Results and Evaluations
-We train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell.
-The training dataset contains 27136 crystal structures.
-The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category.
-On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalance distribution of crystals over space groups in Nature [30].
-In fact, there is no training data in 61 out of 230 space groups as shown in Fig.
-6.
-Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature.
-Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters.
-Next, we will select the model checkpoint with the lowest total validation loss and examine its learned features.
-Then, we will use the model to generate 1000 crystal samples for each of the 230 space groups and evaluate their validity, stability, and novelty.
-These results highlight the benefits of built-in space group symmetries in the model and point to possible future applications of the CrystalFormer model.
-III.1 Learned features
+In the present paper, we will generate crystals using temperature unless mentioned explicitly.
+Besides autoregressive sampling, one can also perform Markov chain Monte Carlo (MCMC) sampling based on the likelihood Eq.
+(1) of the CrystalFormer.
+MCMC sampling of crystals is a way to walk through the crystalline materials starting from an existing crystal structure.
+At each step of the random walk, one proposes an element substitution, atom position shift, or lattice deformation to change the crystal from to , then accepts or rejects the proposal according to the model probability according to the Metropolis acceptance rate .
+MCMC sampling is particularly useful for incorporating additional constraints or guidance in the sampling procedure.
+Moreover, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
+III CrystalFormer learns chemical intuition by compressing materials database
+Nature tends to favor symmetrical crystal structures.
+Crystallographic space groups capture and quantify this inductive bias of nature, thereby significantly simplifying the spaces of crystal materials.
+In light of the space group symmetries, crystals also have an unexpected yet natural sequential and discrete representation, which derives from two tables in nature: the periodic table of elements determined by quantum mechanics and the table of Wyckoff positions of the 230 space groups determined by group theory.
+To construct a certain crystal, we only need to select atoms from the periodic table and place them sequentially into the Wyckoff positions in the unit cell.
+In this crystal language, the “word order” is determined by the alphabetical order of Wyckoff letters, the “grammar” corresponds to the solid-state chemistry rules, and the “synonyms” represent interchangeable elements (Sec.
+III.1), the “sentence length” correspond to atom number in the unit cell (Sec.
+III.2), and the “idioms” correspond to common chemical coordination (Sec.
+III.3).
+CrystalFormer employs an autoregressive transformer to learn the crystal language, thereby exploring yet-to-be-discovered crystalline materials.
+It compresses and internalizes the crystal materials database, expressing solid-state chemical knowledge through neural network parameters; reflecting the associative ability of material space through neural network activations; and describing chemical intuition through the model probability (Sec.
+III.4).
+Similar to generative models used for generating text, images, and videos, CrystalFormer can directly generate “realistic” crystal materials.
+However, rather than worrying about the fake contents of AI-generated media, these AI-generated crystal materials could potentially be synthesized and be useful to human civilization.
+Next, we will inspect the learned features and sample statistics of the model to build up an understanding of the CrystalFormer.
+We carry out inspectations for a few selected space groups.
+The findings are neverthelss general.
+These findings not only provide confidence a bit of understandings of the model, but also direct us to the suitable applications of CrystalFormer.
+III.1 Atom embeddings and chemical similarity
 Figure 3 visualizes the cosine similarity of the learned vector embedding of the chemical species.
 Red colors in the figure indicate similar chemical species identified by the model.
-One clearly sees the periodicity shows up as off-diagonal stripes.
+One sees the chemical similarity within groups of elements show up as off-diagonal red stripes.
 Moreover, there are visible clusters for Lanthanide elements (La-Lu).
 The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc).
-Being able to learn chemical similarities from data [42, 43, 44, 45, 46, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures.
-In particular, the features shown in Fig.
-3 is strikingly similar to the similarity map constructed purposely based on substitution pattern [42] which was later used for substitution-based material discovery [47].
+The features shown in Fig.
+3 are strikingly similar to the similarity map constructed purposely based on substitution pattern [46, 47] which was later used for substitution-based material discoveries [48, 5].
+In the context of language modeling, the chemical similarities correspond to synonyms of chemical species tokens.
+Having the ability to learn them from data [46, 47, 49, 50, 51, 52, 53, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures.
+III.2 Atom number distributions
+The number of atoms corresponds to the length of non-padding atoms in the sequential representation, which is captured well by CrystalFormer.
 Figure 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups.
@@
 Therefore, the space group symmetry is a key control variable for the atom number distribution.
-Building Wyckoff positions information into the CrystalFormer model architecture removes the necessities of querying the training data to find out the number of atoms for a targeted space group [16].
+Incorporating Wyckoff positions information into the CrystalFormer model architecture removes the necessity of querying the training data to find out the number of atoms for a targeted space group [16].
+Recently, Ref.
+[54] reports an abundance of inorganic compounds whose primitive unit cell contains a number of atoms that is a multiple of four.
+There are different ways to reason about the observed ”rule of four” depending on one’s view of how a crystal is formed.
+For example, one can often break inorganic solids into polyhedra as building blocks.
+Otherwise, Ref.
+[55] considers the most probable values of the number of atoms in a formula unit and the number of formula units per primitive cell.
+In line with the discussion here, the ”rule of four” is the combination of three factors 1) the distribution of crystalline materials among space groups [30]; 2) the distribution of atoms in Wyckoff positions [33] of a given space group; and 3) the multiplicities of Wyckoff positions and multiplicities of conventional versus cells.
+The first two are statistical rules determined by the inter-atomic interactions while the third one is a mathematical fact of space group theory.
+In the end, the point we want to make is that the design of CrystalFormer and its associated crystal representation allow it to learn the ”rule of four” and many other to-be-discovered ”rules”, which manifest themselves as marginal statistics of learn probability distribution.
+Most importantly, CrystalFormer will utilize these ”empirical rules” when generating novel yet reasonable crystal samples.
+III.3 Wyckoff-Atom gram
 Figure 5 shows heat maps of Wyckoff positions and chemical species for the space group (No.
 225).
-First of all, one sees that most atoms occupy those special Wyckoff positions with high symmetries.
-Moreover, the heat map reveals interesting solid-state chemistry knowledge as it suggests where each atom tends to appear in a crystal.
-For example, there are vertical stripes at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials.
-Moreover, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities.
+The heat map is analogous to bigram frequency statistics in language modeling.
+In the present context, it reveals interesting solid-state chemistry knowledge related to where each atom tends to appear in a unit cell.
+First of all, one sees that most atoms occupy special Wyckoff positions (Wyckoff letters at the beginning of the alphabet) with higher site symmetries.
+The distribution of generated data is in agreement with test data and recent statistics [33].
+Moreover, there are vertical blanks at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials.
+Lastly, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities [30].
 Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials.
-On the other hand, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset.
+Nevertheless, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset.
 We believe that is due to that the hydrogen element takes only about in the training data for the space group.
 Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
-There has been a long history of summarizing empirical or statistical chemistry rules encoded in materials data and then use them to instruct search of crystal structures [48, 49, 50, 51, 42, 52].
-Our analysis shows that CrystalFormer also ingests a number of chemical intuitions in the training data for generative modelling.
-Since training procedure compresses the chemistry knowledge into a neural network model with tractable likelihoods, it is possible to employ CrystalFormer in the Monte Carlo or evolution strategy search of crystal structures, besides as a probabilistic generative model.
-III.2 Validity and novelty of generated samples
-Figure 6 illustrates the structure and compositional validity of generated samples across all 230 space groups.
-Following the Ref.
-[53], a structure meets the validity criteria if the shortest atomic distance exceeds , a lenient standard.
-Composition validity requires charge neutrality as computed by SMACT [54].
-This is, however, an overly stringent criterion since the composition validity of the training set is only around by this measure [11].
-Note that the CrystalFormer is able to generate reasonable samples even for those space groups without any training data.
-| Space group | Crystal system | Training samples | Validity () | |
-| Struc. | Comp. | |||
-| 2 | Triclinic | 676 | 83.10 | 83.0 |
-| 12 | Monoclinic | 1273 | 87.70 | 81.80 |
-| 62 | Orthorhombic | 1187 | 95.50 | 87.20 |
-| 139 | Tetragonal | 1233 | 97.70 | 83.40 |
-| 166 | Trigonal | 1076 | 98.50 | 85.0 |
-| 194 | Hexagonal | 1129 | 99.40 | 89.90 |
-| 225 | Cubic | 3960 | 99.60 | 93.50 |
-| 1 | Triclinic | 27136 | 91.40 | 80.20 |
-| Autoregressive models | ||||
-| PGSchNet [55] | 99.65 | 75.96 | ||
-| LM-CH (character-level tokenization) [18] | 84.81 | 83.55 | ||
-| LM-AC (atom coordinate-level tokenization) [18] | 95.81 | 88.87 | ||
-| Crystal-LLM [20] | 96.5 | 86.3 | ||
-| Diffusion models | ||||
-| CDVAE [11] | 100.0 | 86.70 | ||
-| DiffCSP [13] | 100.0 | 83.25 | ||
-| DiffCSP++ [26] | 99.94 | 85.12 | ||
-| UniMat-Large [15] | 97.2 | 89.4 | ||
-Table 2 reports the validity of generated samples for selected space groups in each of the seven crystal systems.
-To ensure the numbers are representative, we chose the space group to be the one with the most training data for each crystal system.
-One sees that the model performs better for more symmetric space groups.
-This is a nice feature that complements existing crystal generative models, which mostly have difficulties in generating highly symmetric structures.
-As a reference, we also list the validity of the generated samples suppose one treats the crystals as if they are all in the space group (No.
-1) with only translational symmetry.
-One sees the structure validity scores of space group improves compared to the one shown in Figure 6 due to increased training samples.
-The second part of Table 2 shows reference results in the literature for the same validity test.
-In principle, the performance of the present model should fall back to the language model approaches [18, 20].
-The remaining gap may be due to details such as the including the header line in the crystallographic information files (CIF), specific sampling strategy of language models, or the additional post-selection of samples [20].
-In the table, the DiffCSP++ [26] is the only alternative model that exploits the space group symmetry in the generation process.
-However, different from us, the DiffCSP++ model does not predict Wycokff position as a part of generation process.
-Therefore, DiffCSP++ needs to search for template structure in the training set for generation, which may limit its generality.
-Besides works listed in Table 2 that reported validity scores in a comparable settings, Ref.
-[19] has conditioned the generation of CIF on the space group symbols in a language model setting.
-Ref.
-[16] considered space group conditioned crystal generation using a fine-tuned generative model with space group labels.
-Neither approach provides exact constraints on the space group, which could yield problematic structures for large systems and highly symmetric space groups.
-Ref.
-[23] considered generating symmetric crystals in their Wyckoff representation.
-However, the model does not consider fractional coordinates and lattice parameters, so it requires a subsequent computational search to completely determine the crystal structure.
-Sec.
-IV.2 contains a more in-depth discussion of the relationship of the present approach with related works.
-Figure 7(a) shows the validity of generated samples as a function of sampling temperature.
-One clearly sees that reduced temperature increases the validity of samples at the cost of reducing the diversity [20].
-Figure 7(b) shows the novelty and uniqueness evaluated on generated samples with temperature.
-Novelty quantifies the proportion of new structures in the generated samples that were unseen in the training dataset.
-Uniqueness represents the percentage of distinct, non-redundant structures among the generated samples [56].
-One sees that across different temperatures the uniqueness remains high, indicating the model does not collapse to a mode that produces duplicated samples.
-On the other hand, the model produces close to novel material as temperature increases, which nicely demonstrates modal covering behavior of the maximum likelihood estimation training [57].
-Having a model distribution broader than the span of the dataset is crucial for material discovery.
-III.3 Stability of generative samples
-Figure 8(a) shows the histogram of the log-likelihood of generated and test samples show nice agreement.
-We also visualize structures of a few generated samples which deems to be very likely, typical, and unlikely according to their log-likelihoods.
-Interestingly, we did not observe a clear correlation between the model likelihood and energy of the structure.
-Fig.
-8(b) shows the energy above hull based on energies calculated with density functional theory (DFT).
-One sees the distributions of unrelaxed and relaxed structures are very similar, suggesting that the generated structures are already very close to local minima of DFT calculations.
-For the given computational budget detailed in the Appendix B, we manage to reach convergence for 973 out of 1000 structures in the DFT relaxation.
-All of the structures retains the same space group symmetry after the DFT relaxation 222Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No.
-221)..
-With its default parameters, StructureMatcher from pymatgen [60] is able to match all 973 structures before and after relaxation.
-The average root mean squared displacement (RMSD) normalized by computed for these matched structures is 0.0061, which confirms the generated samples are indeed very close to DFT local minima.
-Figure 8(b) shows a pronounced peak around eV/atom showing the model indeed generates a large fraction of candidates of metastable materials [58].
-Finally, we found 21 relaxed samples are stable by the criteria and summarize them in Table S2 of appendix B.
-Most of the found stable materials only have atoms occupying the first few Wyckoff positions.
-Among these 21 structures, we found 10 are new crystalline materials that are not contained in the Materials Project database [61].
-IV Discussions
-IV.1 Applications
-With the abilities demonstrated in Sec. III, we believe that CrystalFormer is ready to fit into the existing materials discovery workflow in the following two ways.
-IV.1.1 Initialization for structure search
-Crystal structure search relies heavily on utilizing space group symmetries.
-For example, it is a common practice for crystal structure prediction software [62, 63, 64, 65, 66] and crystal structure search [67, 68] to place atoms at Wyckoff positions to prepare symmetric initial states for subsequent optimization.
-However, such initialization approach faces combinatorial difficulty as the number of chemical species and atoms in the unit cell grow.
-The CrystalFormer is ready to act as a drop-in replacement of random structure initialization.
-In this way, one bypasses the curse-of-dimensionality of exact enumeration with a data-driven probabilistic approach.
-Moreover, the ability of CrystalFormer to generate diverse and close to metastable structures can greatly reduce computational costs of downstream ab initio calculations.
-IV.1.2 Mutation of existing structures
-Mutatation of known crystals is a prominent approach to materials discovery.
-For example, one can employ machine machine-learned force field to relax crystal structures [69, 5, 70, 71] after element substitutions.
+Along the same line of thoughts, coordination polyhedra [56] and lattice structure [57] manifest themselves as higher-order n-gram correlations of Wyckoff position and atom species in the crystal language, which will be captured by the CrystalFormer.
+There has been a long history of mining empirical chemistry rules encoded in materials data and then using them to instruct the search of crystal structures [58, 59, 60, 61, 47, 62].
+Our analysis shows that CrystalFormer ingests chemical intuition, be it speakable or unspeakable, in the training data for generating new materials.
+III.4 Crystal likelihoods
+CrystalFormer compresses chemistry knowledge stored in the material dataset into its parameters.
+In addition to generating crystal samples, CrystalFormer can also compute the likelihoods of crystals via Eq.
+(1).
+Therefore, it is possible to employ CrystalFormer in the Monte Carlo search of crystal structures besides sampling materials directly as a generative model.
+Figure 6 shows the agreement of the likelihoods of generated samples and samples in the test dataset.
+We also visualize structures of a few generated samples which are deemed to be very likely, typical, and unlikely according to their likelihood values.
+We have checked that likelihood is related to the energy of the crystal by locally perturbing the fractional coordinates and lattice parameters.
+However, we did not observe a correlation between the likelihood of these crystals and their energies on a global scale.
+We envision the landscape of likelihoods is much less rough compared to the potential energy surface of crystalline materials.
+Intuitively, it means that the Crystalformer compresses the materials space into a more compact space without many holes that correspond to infeasible high energy states.
+Therefore, likelihood-based exploration of the crystal space discussed in Sec.
+II.3 can be more efficient compared to traditional sampling approaches based on the physical energy functions.
+IV Applications
+We now move on to the practical applications of CrystalFormer to materials discovery and design.
+Compared to many existing materials generation models, CrystalFormer offers precise control over space group symmetry and enables efficient computation of model likelihood.
+These unique features open a wide range of possibilities for integrating it with existing computational software and machine-learning models in a flexible way as we demonstrate below.
+For these applications, we have excluded radioactive elements from the samples [32].
+IV.1 Symmetry-conditioned random structure initialization
+Crystal structure prediction has long been the dream of solid-state chemistry and computational material science researchers [63].
+Typical crystal structure prediction workflow consists of two steps.
+First, one randomly initializes a batch of diverse crystal structures as candidates.
+Second, one optimizes the crystal structures via local and global optimization strategies.
+Utilizing space group symmetries plays a crucial role in both steps, as symmetry enlarges the span of the energy distribution [64, 65, 66] and reduces the search space.
+It is a common practice for crystal structure prediction software [67, 68, 69, 65, 70, 66] and structure search [71, 72, 73] to exploit space group symmetry in the crystal structure initialization.
+However, such an initialization approach faces combinatorial difficulty as the number of chemical species and atoms in the unit cell grows.
+The CrystalFormer is ready to act as a drop-in replacement of random structure initialization for crystal structure prediction.
+In this way, one bypasses the curse-of-dimensionality of exact enumeration [65] with a data-driven probabilistic approach.
+Moreover, the ability of CrystalFormer to generate diverse and near-stable structures can greatly reduce the computational costs of downstream optimizations.
+We select seven space groups (No.
+2), (No.
+12), (No.
+62), (No.
+139), (No.
+166), (No.
+194), and (No.
+225) as representatives of the seven crystal systems.
+We randomly generate 100 crystals for each space group using CrystalFormer.
+On the other hand, we employ PyXtal [66] to generate crystal samples with the same stoichiometry in the same space groups.
+We then carry out structure relaxation using density functional (DFT) calculations.
+| Space group | Crystal system | Converged structures | Retain symmetry | RMSD 1 () | 1 (eV/atom) | ||||
+| CrystalFormer | PyXtal | CrystalFormer | PyXtal | CrystalFormer | PyXtal | CrystalFormer | PyXtal | ||
+| (No. 2) | Triclinic | 46 | 67 | 45 | 67 | 1.929 | 2.038 | 1.034 | 0.913 |
+| (No. 12) | Monoclinic | 55 | 72 | 53 | 67 | 1.846 | 2.235 | 1.233 | 1.660 |
+| (No. 62) | Orthorhombic | 77 | 83 | 76 | 66 | 1.086 | 1.940 | 0.313 | 1.633 |
+| (No. 139) | Tetragonal | 91 | 81 | 88 | 63 | 0.405 | 1.160 | 0.240 | 1.100 |
+| (No. 166) | Trigonal | 83 | 74 | 80 | 71 | 1.028 | 2.183 | 0.352 | 2.489 |
+| (No. 194) | Hexagonal | 97 | 77 | 96 | 60 | 0.355 | 2.129 | 0.324 | 4.100 |
+| (No. 225) | Cubic | 98 | 96 | 95 | 92 | 0.076 | 0.518 | 0.214 | 0.483 |
+| Average | 78.1 | 78.6 | 76.1 | 69.4 | 0.961 | 1.743 | 0.530 | 1.769 | |
+1
+Calculated on the converged structures.
+Figure 7(a)-(g) shows the average energy difference to the energy of final structures versus DFT relaxation steps.
+We neglected the structures whose energy changes and energy change intervals per step during relaxations exceeded 10 eV/atom to eliminate the impact of erroneous steps.
+One sees that CrystalFormer samples generally reach lower energies in fewer relaxation steps.
+This is especially true for space groups with higher symmetries.
+The ability to initialize diverse and high-quality crystal structures enables one to discover more stable materials faster.
+Figure 7(h) shows the histogram of energy above the convex hull constructed by the Materials Project database.
+The dashed line denotes the criterion eV/atom [74] for selecting stable materials.
+Among these candidates, we found 34 and 12 relaxed structures with CrystalFormer and PyXtal initializations that are not contained in the MP-20 dataset.
+We summarize them in Table LABEL:table:samples and Table LABEL:table:pyxtal_samples of appendix C.
+Table LABEL:table:700-stat lists detailed statistics of structure-relaxed samples in seven representative space groups.
+Overall, we find that the CrystalFormer generated structures are of higher quality, especially for those space groups with higher symmetry.
+This observation is supported by the fact that the DFT relaxation often retains the space group symmetry.
+The average root mean squared displacement (RMSD) computed for these converged structures is 0.961 and 1.743 for CrystalFormer and PyXtal initialization, respectively.
+In addition, the average energy above the convex hull also confirms the samples generated by CrystalFormer are indeed much closer to the DFT local minimum than PyXtal initialization.
+IV.2 Structure-conditioned element substitution
+Mutation of known crystals is a prominent approach to materials discovery.
+For example, one can employ a machine-learned force field to relax crystal structures [75, 5, 76, 77] after element substitutions.
 In the lens of generative modeling, the machine learning force field can be regarded as the energy-based model or Boltzmann machines.
 A potential drawback of exploring materials space with an energy-based model is the slow mixing or even ergodicity issue posed by the rough landscape of the potential energy surface.
-In this sense, element substitutions provide a variety of initial seeds, compensating the limitation potential energy surface-based exploration.
+In this sense, element substitutions provide a variety of initial seeds, compensating for the limitation of energy-based exploration.
 Having an alternative measure of crystal likelihood other than the potential energy surface opens a way to employ the model likelihood as a guide for structure search.
-For example, with a trained CrystalFormer, one can perform Markov chain Monte Carlo (MCMC) random walk in the crystalline materials starting from an existing crystal structure.
-At each step of the random walk, one proposes an element substitution, atom shift, or lattice deformation and accept or reject the proposal according to the model likelihood.
-In the long time limit, one will sample from the model distribution.
-However, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
-IV.2 Related works
-Crystal generative models have been explored using variational autoencoder [72, 23], generative adversarial networks [73, 12], normalizing flows [74, 75, 76], diffusion models [11, 12, 13, 26, 14, 15, 16, 21, 22], GFlowNet [24, 25], and autoregressive models [77, 55, 78, 17, 18, 19, 19].
-In these autoregressive models, one either use atomistic features [77, 55, 78] or use pure text tokens [17, 18, 19, 20].
-However, with the introduction of specialized tokens for crystals, the boundary between the two is blurred.
-The CrystalFormer is most closely related to the autoregressvie generative model originally designed for molecules [77, 55, 78].
-However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of atoms in the unit cell.
-Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around rotational equivariance.
+Many crystal structures can be traced back to a few simple, highly symmetrical types.
+Numerous crystals share the same structural prototype but differ in composition, such as perovskite (ABX3), spinel (AB2X4), fluorite (AX2), and so on.
+Figure 8(a) shows double perovskite crystal structures A2BB’X6 which belong to the (No.
+225) space group.
+There are hundreds of known double perovskites with significant interests in their semiconducting, ferroelectric, thermoelectric, and superconducting properties [78].
+Finding more stable materials with this structure prototype using brute force enumeration and high-throughput calculation is a computationally demanding task [79].
+We will generate new double perovskites with CrystalFormer and demonstrate its advantage of over standard element substitution methods.
+Figure 8(a) shows the string representation of double perovskites.
+To generate candidates of double perovskites, we use CrystalFormer to carry out string in-filling tasks.
+Since the autoregressive sampling of the atoms is insufficient to take into account non-causal information in the sequence, we employ MCMC to sweep through the sequence and update chemical species and fractional coordinates [80].
+The acceptance rate for these MCMC updates makes use of the marginalized probability for elements and fractional coordinates as the lattice parameter that appears at the end of the sequence can be integrated.
+Only after the MCMC sampling has been thermalized, we sample the lattice parameters autoregressive to account for the adjustment of the unit cell for given atoms and occupations.
+We use CrystalFormer to generate 100 candidates as the initial DFT relaxation.
+As a comparison, we also employ the SubstitutionPredictorTransformation function [46] implemented in pymatgen [81] to perform element substitution for the crystals with double perovskite structures in the training dataset.
+The substitution probabilities come from data-mining of ICSD dataset [46].
+After the substitution, we use DLSVolumePredictor [82] function of pymatgen to predict the volume of the structure.
+This lattice scaling scheme relies on data-mined bond lengths to predict the crystal volume of a given structure.
+To collect 100 candidates in the ionic substitution approach we have set the probability threshold of SubstitutionPredictorTransformation to 0.01, which is smaller than the typical values adopted in Ref.
+[48].
+The RMSD computed for the DFT-relexed structures is 0.211 and 0.342 for CrystalFormer and ionic substitution [46], respectively.
+This observation confirms that the structures generated by CrystalFormer are closer to the DFT local minimum than ionic substitution [46].
+Moreover, Figure 8(b) shows the histogram of energy above the convex hull of the Materials Project database.
+Overall, CrystalFormer and ionic substitution [46] found 9 and 3 double perovskites with eV/atom which are not contained in the MP-20 dataset, details in appendix D.
+The superior performance of CrystalFormer-guided MCMC is understandable since its likelihood takes into account the context of space group and atomic environment rather than marginal two-body correlation [46] in ionic substitution.
+The ionic substitution approach also shows two limitations in practical applications.
+First, some of the ions in the compound can not be substituted as they are missing in the probability table.
+Second, the approach relies on the calculability of the elements’ valence states.
+As a final remark, although the discussion here focuses on generating crystals with given prototype structures, the generation of crystals with a given crystal lattice [83] is also feasible with CrystalFormer.
+This is because the crystal lattice can be straightforwardly expressed as constraints on the space group and occupied Wyckoff letters [57].
+IV.3 Plug-and-play materials design
+Finally, we demonstrate CrystalFormer’s ability to aid property-guided exploration of crystalline materials in a versatile and flexible manner.
+The trained CrystalFormer captures the space group conditioned crystal probability , which we treat as a prior probability for stable crystals.
+By combining it with a crystal property prediction model that provides the forward likelihood probability , one can carry out property-guided materials generation in a plug-and-play manner.
+According to Bayes’ rule, the posterior for crystals given property reads
+By sampling from this posterior distribution, one can generate crystal samples with property guidance.
+Since the posterior probability Eq.
+(3) typically does not process autoregressive property with respect to , we carry out MCMC sampling to sample from the posterior distribution [84].
+The plug-and-play feature makes designing crystalline materials in this way particularly appealing because it is possible to apply multiple conditions by simply adding log-likelihoods from multiple predictors.
+The framework applies to the inverse problem of solving cyrystal structures based on experimently observed diffraction spectra equally well [85, 86], where the goal is to simultaneously optimizing the matching probability to experimental observation and stability of the crystal.
+Any property prediction model can used in conjunction with CrystalFormer for property-guided material generation.
+We train two crystal property prediction models using the MP-20 dataset and achieve mean absolute error (MAE) of 0.325 eV in band gap and MAE of 0.110 eV/atom in formation energy, which is comparable to the crystal graph convolutional neural network [87] trained on the same dataset [88].
+We use the output of these two property prediction models as the forward probability of Eq.
+(3).
+More details are in appendix E.
+Figure 9(a) demonstrates the controlled generation of materials with target band gap at and the formation of energy crystals [89, 90].
+The conditional probability Eq.
+(3) contains both the model likelihood and the property regression MAE.
+Therefore, the generated samples will strike a balance between the two.
+To draw samples from the conditional probability distribution, we randomly generate a batch of 1000 crystal samples and sweep through the crystal sequence to update the atom species, fractional coordinates, and lattice parameters 222For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling..
+Monte Carlo update of chemical species so to achieve the desired properties can be regarded as a systematic data-informed way of carrying out cation-transmutation for materials inverse design [92].
+After reaching equilibrium, the histogram of band gap and formation energy is centered around the target values, which is shifted significantly away from the value of unconditionally generated samples.
+On the other hand, the likelihood shown in Fig.
+9(b) indicates these conditional-generated crystals are not typical samples with respect to the unconditional distributuion.
+Neverthless, they are still probable samples according to the crystal prior given by the CrystalFormer.
+Given a myriad of materials property prediction models developed over the years and the inconvenience of re-training or fine-tuning the foundational generative model [16], we envision the plug-and-play generation approach demonstrated here to be a scalable way for materials design.
+We have exposed an interface of CrystalFormer in the code repository [34] for users to plug in arbitrary conditioners for guided materials generation.
+V Related works
+Crystal generative models have been explored using variational autoencoder [93, 23], generative adversarial networks [94, 12], normalizing flows [95, 96, 97, 98], diffusion models [11, 12, 13, 26, 14, 15, 16, 21, 22], GFlowNet [24, 25], and autoregressive models [99, 100, 101, 17, 18, 19, 20].
+In these autoregressive models, one either uses atomistic features [99, 100, 101, 17] or uses pure text descriptions [18, 19, 20].
+Nevertheless, with the introduction of specialized tokens for crystals, the boundary between the two is blurred.
+The CrystalFormer is most closely related to the autoregressive generative model originally designed for molecules [99, 100, 101].
+However, instead of predicting the relative distances of atoms, we predict the Wyckoff positions of symmetry-inequivalent atoms in the unit cell.
+Having the luxury of the space group symmetry for crystals provides strong hints on where to put the atoms in the unit cell and greatly simplifies the design around spatial symmetries.
 On the other hand, compared to Refs.
-[17, 18, 19, 20] which treat text descriptions of crystals using autoregressive language models, we are only dealing with more concise and essential atomistic representation of crystals, which leads to smaller model size and faster sampling speed.
-Fast generation speed is not only a welcoming feature but also will be crucial for intelligent exploration of materials space based on combinations of probabilistic generation and post-selection, backtracking, and searching techniques [79].
-More importantly, CrystalFormer guarantees space group constraints since these symmetries are baked in the model architecture rather than learned as statistical correlation from text data.
+[19] which treats text descriptions of crystals using autoregressive language models, we are only dealing with a more concise and essential atomistic representation of crystals, which leads to a smaller model size and faster sampling speed.
+Fast generation speed is not only a welcoming feature but also will be crucial for further exploration of materials space based on combinations of probabilistic generation and post-selection, Monte Carlo sampling, backtracking, and searching techniques [102].
+More importantly, by baking in the space group symmetry in the model rather than learning them as statistical correlation from texts [18, 20], CrystalFormer guarantees space group constraints and cherishes the precious data and computing time.
 In a sense, the present work employs intrinsic mathematical (as opposed to natural) language to incorporate the symmetry principle in the generative modeling of crystals.
 As a side remark, the Wyckoff position features have been used in machine learning models for materials property prediction [49, 103].
-Incorporating space group information in the encoder-only transformer models may also boost their property prediction performance [82, 83] as suggested by the related study [84].
-V Outlook
-Crystal structure prediction has long been the dream of computational material sciences [85].
-CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one framework.
+Incorporating space group information in the encoder-only transformer models may also enhance their property prediction performance [104, 105, 106] as suggested by Ref.
+[107].
+VI Outlook
+Precisely controlling the space group in the generative model of crystalline materials not only greatly simplifies the task but also is a highly desired feature for materials discovery and design.
+CrystalFormer integrates exact symmetry principles from math and empirical chemical intuitions from data into one unified framework.
 Probabilistic generative modeling of crystalline materials using CrystalFormer opens the way to many future innovations in materials discovery.
-Precisely controlling the space group in the generative model of crystalline materials is not only a highly desired feature but also greatly simplifies the task.
+Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16].
 An obvious future direction is to scale up the model as well as the training dataset, especially curating a dataset with better coverage of space groups.
-Note that the MP-20 dataset has by no means exhausted all available crystalline material [19, 16].
-The transformer-based generative model is ready to be scaled up with more training data, in the same fashion as large language models [86].
+In particular, extending the dataset to include both inorganic and organic crystals [108] may be beneficial as it improves the data coverage of low symmetric space groups.
+The transformer-based generative model is ready to be scaled up to work with much larger and more diverse training data, in the same fashion as large language models [109].
 Given similar model architectures, the idea of generative pretraining of a foundational model for material generation is appealing.
 When scaling up the model it will be interesting to note the possible appearance of neural scaling law [110] as it has also been showing up in other contexts of atomistic modeling [111].
-The model architecture is still open to further improvement to better serve the purpose of material discovery, First of all, to better facilitate data efficiency learning and structure phase transitions-related applications, it will be useful to further exploit the group-subgroup structure [89] in the model architecture.
-Second, besides chemical elements, it will be useful to constrain the stoichiometry of generated materials.
-This may be achieved by tuning the logit bias or combining the present probabilistic generation approaches with traditional enumeration approaches [90].
-Lastly, it is worth exploring to use CrystalFormer as the base distribution in the flow model and employ symmetry-persevering transportation’s to further adjust the atoms coordinates and unit cells [10, 26], which mimics a symmetry-constrained relaxation process [91].
-Conditioned crystal structure generation based on material properties [78, 16] and experimental measurements [92] are both straightforward with CrystalFormer.
-One can either extend the space group embedding or employ the full encoder-decoder transformer architecture [36] for conditioned crystal generation.
-The first step of property-guided generation could take into account of synthesizability and stability of materials.
-In the future, proposing materials with multiple properties is a promising direction for material generative models.
-To achieve that goal, curating a high-quality material dataset with either computed or measured properties is crucial.
-Needless to say, to close the loop of material discovery, one will need to carry out experimental verification for the generated materials.
+The model architecture and sampling strategy are both open to further refinement to better serve the purpose of material discovery.
+First of all, to better facilitate data efficiency learning and structure phase transitions-related applications, it will be useful to further exploit the Euclidean normalizer [112] and group-subgroup relation [113] in the model architecture or training procedure.
+Second, it is worth exploring using CrystalFormer as the base distribution in the flow model and employs symmetry-persevering transportation to further adjust the atoms coordinates and unit cells [10, 26], which mimics a symmetry-constrained relaxation process [114].
+Lastly, it may be worth employing more advanced constrained and guided sequence generation methods [115, 116, 117, 118] for more flexible control on the elements, structure, or stoichiometry of generated materials.
+Conditioned materials generation depending on properties [101, 16, 21, 22] and experimental measurements [119] are highly desired features of materials generative model.
+Although it is straightforward to extend CrystalFormer (e.g.
+extend the space group embedding or employ the encoder-decoder transformer architecture [40]) to incorporate these conditions, we are particularly excited about the plug-and-play routine demonstrated in Sec.
+IV.3.
+Along this line, we envision an ecosystem [120] where the foundational generative model for and more specialized discriminative models are developed separately but brought together via the Bayes rule.
 Acknowledgements.
-We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Qi Yang, Shigang Ou, Xinyang Dong, and Quansheng Wu for useful discussions.
+We thank Han Wang, Lin Yao, Linfeng Zhang, Chen Fang, Yanchao Wang, Zhenyu Wang, Qi Yang, Shigang Ou, Xinyang Dong, Wenbing Huang, Quansheng Wu, Wanjian Yin, Xi Dai, Shuang Jia, Hangtian Zhu, Jiangang Guo, and Hongjian Zhao for useful discussions.
 This project is supported by the National Natural Science Foundation of China under Grants No.
@@
 - Luo et al. [2024] X. Luo, Z. Wang, P. Gao, J. Lv, Y. Wang, C. Chen, and Y. Ma, Deep learning generative model for crystal structure prediction, (2024), arXiv:2403.10846 [cond-mat.mtrl-sci] .
-- Ye et al. [2024] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, (2024), arXiv:2403.12478 [cond-mat.mtrl-sci] .
-- Zhu et al. [2023] R. Zhu, W. Nong, S. Yamazaki, and K. Hippalgaonkar, Wycryst: Wyckoff inorganic crystal generator framework, (2023), arXiv:2311.17916 [cond-mat.mtrl-sci] .
+- Ye et al. [2024a] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, (2024a), arXiv:2403.12478 [cond-mat.mtrl-sci] .
+- Zhu et al. [2024] R. Zhu, W. Nong, S. Yamazaki, and K. Hippalgaonkar, Wycryst: Wyckoff inorganic crystal generator framework, Matter doi.org/10.1016/j.matt.2024.05.042 (2024).
 - AI4Science et al.
@@
 - Marsh [1999] R. E. Marsh, P1 or P1¯? Or something else?, Acta Crystallographica Section B 55, 931 (1999).
+- Cheetham and Seshadri [2024] A.
+K.
+Cheetham and R.
+Seshadri, Artificial intelligence driving materials discovery?
+perspective on the article: Scaling deep learning for materials discovery, Chemistry of Materials 36, 3490 (2024).
+- Hornfeck [2022] W. Hornfeck, On the combinatorics of crystal structures: number of wyckoff sequences of given length, Acta Crystallographica Section A: Foundations and Advances 78, 149 (2022).
 - [34] See https://github.com/deepmodeling/CrystalFormer for code and model checkpoint.
@@
 Hu, Machine learning-based prediction of crystal systems and space groups from inorganic materials compositions, ACS Omega 5, 3596 (2020), pMID: 32118175.
+- Liang et al. [2020] H. Liang, V. Stanev, A. G. Kusne, and I. Takeuchi, Cryspnet: Crystal structure predictions via neural networks, Phys. Rev. Mater. 4, 123802 (2020).
 - Wang et al. [2023a] D.-Y. Wang, H.-F. Lv, and X.-J. Wu, Crystallographic groups prediction from chemical composition via deep learning, Chinese Journal of Chemical Physics 36, 66 (2023a).
+- Venkatraman and Carvalho [2024] V. Venkatraman and P. A. Carvalho, Accurate space-group prediction from composition, Journal of Applied Crystallography (2024).
 - Hahn et al. [1983] T. Hahn, U. Shmueli, and J. W. Arthur, International tables for crystallography, Vol. 1 (Reidel Dordrecht, 1983).
@@
 - Note [1] We follow the convention that capital letters appear after lower case letters.
-This handles the edge case of the No.
-47 space group whose Wyckoff positions reach ’A’.
-This rule also demands that the padded atoms appear at the end of the sequence with Wyckoff letter ’X’.
+This handles the edge case of the space group (No.
+47) whose Wyckoff letters used up 26 lowercase letters and reached ’A’ for the generic position.
+In addition, we use the letter ’X’ to indicate the Wyckoff position of padding atoms that appear at the end of the sequence, see e.g.
+of Fig.
+1(b).
 - [43] OpenAI, Using logit bias to alter token probability with the openai api, OpenAI Help Center, retrieved March 14, 2024.
@@
 - Parthé and Gelato [1984] E. Parthé and L. M. Gelato, The standardization of inorganic crystal-structure data, Acta Crystallographica Section A 40, 169 (1984).
+- Hautier et al. [2011] G. Hautier, C. Fischer, V. Ehrlacher, A. Jain, and G. Ceder, Data mined ionic substitutions for the discovery of new compounds, Inorganic chemistry 50, 656 (2011).
 - Glawe et al.
@@
 Marques, The optimal one dimensional periodic table: a modified pettifor chemical scale from data mining, New Journal of Physics 18, 093011 (2016).
+- Wang et al. [2021] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021).
+- Jain and Bligaard [2018] A. Jain and T. Bligaard, Atomic-position independent descriptor for machine learning of material properties, Phys. Rev. B 98, 214112 (2018).
 - Zhou et al. [2018] Q. Zhou, P. Tang, S. Liu, J. Pan, Q. Yan, and S.-C. Zhang, Learning atoms for materials discovery, Proceedings of the National Academy of Sciences 115, E6411 (2018).
@@
 Gurlo, Crabnet for explainable deep learning in materials science: bridging the gap between academia and industry, Integrating Materials and Manufacturing Innovation 11, 41 (2022).
-- Wang et al. [2021a] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021a).
+- Gazzarrini et al.
+[2024] E.
+Gazzarrini, R.
+K.
+Cersonsky, M.
+Bercx, C.
+S.
+Adorf, and N.
+Marzari, The rule of four: anomalous distributions in the stoichiometries of inorganic compounds, Npj Computational Materials 10, 73 (2024).
+- Palgrave [2024] R. Palgrave, An explanation for the rule of four in inorganic materials 10.26434/chemrxiv-2024-sxqwh (2024).
+- Alvarez [2005] S. Alvarez, Polyhedra in (inorganic) chemistry, Dalton Transactions , 2209 (2005).
+- Regnault et al.
+[2022] N.
+Regnault, Y.
+Xu, M.-R.
+Li, D.-S.
+Ma, M.
+Jovanovic, A.
+Yazdani, S.
+S.
+Parkin, C.
+Felser, L.
+M.
+Schoop, N.
+P.
+Ong, et al., Catalogue of flat-band stoichiometric materials, Nature 603, 824 (2022).
 - Pauling [1929] L. Pauling, The principles determining the structure of complex ionic crystals, Journal of the American Chemical Society 51, 1010 (1929).
@@
 - Allahyari and Oganov [2020] Z. Allahyari and A. R. Oganov, Coevolutionary search for optimal materials in the space of all possible compounds, npj Computational Materials 6, 55 (2020).
-- Court et al.
-[2020] C.
-J.
-Court, B.
-Yildirim, A.
-Jain, and J.
-M.
-Cole, 3-d inorganic crystal structure generation and property prediction via representation learning, Journal of Chemical Information and Modeling 60, 4518 (2020), pMID: 32866381.
-- Davies et al.
-[2019] D.
-W.
-Davies, K.
-T.
-Butler, A.
-J.
-Jackson, J.
-M.
-Skelton, K.
-Morita, and A.
-Walsh, Smact: Semiconducting materials by analogy and chemical theory, Journal of Open Source Software 4, 1361 (2019).
-- Gebauer et al.
-[2019] N.
-Gebauer, M.
-Gastegger, and K.
-Schütt, Symmetry-adapted generation of 3d point sets for the targeted discovery of molecules, in Advances in Neural Information Processing Systems, Vol.
-32 (Curran Associates, Inc., 2019).
-- [56] https://github.com/sparks-baird/matbench-genmetrics.
-- Goodfellow et al. [2016] I. Goodfellow, Y. Bengio, and A. Courville, Deep Learning (MIT Press, 2016) http://www.deeplearningbook.org.
+- Maddox [1988] J. Maddox, Crystals from first principles, Nature 335, 201 (1988).
+- Wales [1998] D. J. Wales, Symmetry, near-symmetry and energetics, Chemical physics letters 285, 330 (1998).
+- Avery and Zurek [2017] P. Avery and E. Zurek, Randspg: An open-source program for generating atomistic crystal structures with specific spacegroups, Computer Physics Communications 213, 208 (2017).
+- Fredericks et al.
+[2021] S.
+Fredericks, K.
+Parrish, D.
+Sayre, and Q.
+Zhu, Pyxtal: A python library for crystal structure generation and symmetry analysis, Computer Physics Communications 261, 107810 (2021).
+- Wang et al. [2010] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Crystal structure prediction via particle-swarm optimization, Phys. Rev. B 82, 094116 (2010).
+- Pickard and Needs [2011] C. J. Pickard and R. J. Needs, Ab initio random structure searching, Journal of Physics: Condensed Matter 23, 053201 (2011).
+- Lyakhov et al. [2013] A. O. Lyakhov, A. R. Oganov, H. T. Stokes, and Q. Zhu, New developments in evolutionary structure prediction algorithm uspex, Computer Physics Communications 184, 1172 (2013).
+- Falls et al.
+[2020] Z.
+Falls, P.
+Avery, X.
+Wang, K.
+P.
+Hilleke, and E.
+Zurek, The xtalopt evolutionary algorithm for crystal structure prediction, The Journal of Physical Chemistry C 125, 1601 (2020).
+- Cheng et al. [2022] G. Cheng, X.-G. Gong, and W.-J. Yin, Crystal structure prediction by combining graph network and optimization algorithm, Nature Communications 13, 1492 (2022).
+- Wang et al.
+[2023b] H.-C.
+Wang, J.
+Schmidt, M.
+A.
+L.
+Marques, L.
+Wirtz, and A.
+H.
+Romero, Symmetry-based computational search for novel binary and ternary 2d materials, 2D Materials 10, 035007 (2023b).
+- Zhang et al.
+[2024] Q.
+Zhang, A.
+Choudhury, and A.
+Chernatynskiy, A symmetry-oriented crystal structure prediction method for crystals with rigid bodies, (2024), arXiv:2407.21337 [cond-mat.mtrl-sci] .
 - Sun et al.
@@
 Ceder, The thermodynamic scale of inorganic crystalline metastability, Science Advances 2, e1600225 (2016).
-- Note [2] Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No. 221).
-- Ong et al.
-[2013] S.
-P.
-Ong, W.
-D.
-Richards, A.
-Jain, G.
-Hautier, M.
-Kocher, S.
-Cholia, D.
-Gunter, V.
-L.
-Chevrier, K.
-A.
-Persson, and G.
-Ceder, Python materials genomics (pymatgen): A robust, open-source python library for materials analysis, Computational Materials Science 68, 314 (2013).
-- Jain et al.
-[2013] A.
-Jain, S.
-P.
-Ong, G.
-Hautier, W.
-Chen, W.
-D.
-Richards, S.
-Dacek, S.
-Cholia, D.
-Gunter, D.
-Skinner, G.
-Ceder, and K.
-A.
-Persson, Commentary: The Materials Project: A materials genome approach to accelerating materials innovation, APL Materials 1, 011002 (2013).
-- Pickard and Needs [2011] C. J. Pickard and R. J. Needs, Ab initio random structure searching, Journal of Physics: Condensed Matter 23, 053201 (2011).
-- Wang et al. [2012] Y. Wang, J. Lv, L. Zhu, and Y. Ma, Calypso: A method for crystal structure prediction, Computer Physics Communications 183, 2063 (2012).
-- Lyakhov et al. [2013] A. O. Lyakhov, A. R. Oganov, H. T. Stokes, and Q. Zhu, New developments in evolutionary structure prediction algorithm uspex, Computer Physics Communications 184, 1172 (2013).
-- Avery and Zurek [2017] P. Avery and E. Zurek, Randspg: An open-source program for generating atomistic crystal structures with specific spacegroups, Computer Physics Communications 213, 208 (2017).
-- Fredericks et al.
-[2021] S.
-Fredericks, K.
-Parrish, D.
-Sayre, and Q.
-Zhu, Pyxtal: A python library for crystal structure generation and symmetry analysis, Computer Physics Communications 261, 107810 (2021).
-- Cheng et al. [2022] G. Cheng, X.-G. Gong, and W.-J. Yin, Crystal structure prediction by combining graph network and optimization algorithm, Nature Communications 13, 1492 (2022).
-- Wang et al.
-[2023b] H.-C.
-Wang, J.
-Schmidt, M.
-A.
-L.
-Marques, L.
-Wirtz, and A.
-H.
-Romero, Symmetry-based computational search for novel binary and ternary 2d materials, 2D Materials 10, 035007 (2023b).
 - Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
@@
 Csányi, A foundation model for atomistic materials chemistry, (2024), arXiv:2401.00096 [physics.chem-ph] .
+- Vasala and Karppinen [2015] S. Vasala and M. Karppinen, A2b’b”o6 perovskites: a review, Progress in solid state chemistry 43, 1 (2015).
+- Wang et al. [2024] Y. Wang, B. Baldassarri, J. Shen, J. He, and C. Wolverton, Landscape of thermodynamic stabilities of a2bb’o6 compounds, Chemistry of Materials (2024).
+- Miao et al.
+[2019] N.
+Miao, H.
+Zhou, L.
+Mou, R.
+Yan, and L.
+Li, CGMH: Constrained sentence generation by metropolis-hastings sampling, in Proceedings of the AAAI Conference on Artificial Intelligence, Vol.
+33 (2019) pp.
+6834–6842.
+- Ong et al.
+[2013] S.
+P.
+Ong, W.
+D.
+Richards, A.
+Jain, G.
+Hautier, M.
+Kocher, S.
+Cholia, D.
+Gunter, V.
+L.
+Chevrier, K.
+A.
+Persson, and G.
+Ceder, Python materials genomics (pymatgen): A robust, open-source python library for materials analysis, Computational Materials Science 68, 314 (2013).
+- Chu et al. [2018] I.-H. Chu, S. Roychowdhury, D. Han, A. Jain, and S. P. Ong, Predicting the volumes of crystals, Computational Materials Science 146, 184 (2018).
+- Okabe et al.
+[2024] R.
+Okabe, M.
+Cheng, A.
+Chotrattanapituk, N.
+T.
+Hung, X.
+Fu, B.
+Han, Y.
+Wang, W.
+Xie, R.
+J.
+Cava, T.
+S.
+Jaakkola, Y.
+Cheng, and M.
+Li, Structural constraint integration in generative model for discovery of quantum material candidates, (2024), arXiv:2407.04557 [cond-mat.mtrl-sci] .
+- Verkuil et al.
+[2022] R.
+Verkuil, O.
+Kabeli, Y.
+Du, B.
+I.
+Wicky, L.
+F.
+Milles, J.
+Dauparas, D.
+Baker, S.
+Ovchinnikov, T.
+Sercu, and A.
+Rives, Language models generalize beyond natural proteins, BioRxiv , 2022 (2022).
+- Meredig and Wolverton [2013] B. Meredig and C. Wolverton, A hybrid computational–experimental approach for automated crystal structure solution, Nature materials 12, 123 (2013).
+- Parackal et al.
+[2023] A.
+S.
+Parackal, R.
+E.
+A.
+Goodall, F.
+A.
+Faber, and R.
+Armiento, Identifying crystal structures beyond known prototypes from x-ray powder diffraction spectra, (2023), arXiv:2309.16454 [cond-mat.mtrl-sci] .
+- Xie and Grossman [2018] T.
+Xie and J.
+C.
+Grossman, Crystal graph convolutional neural networks for an accurate and interpretable prediction of material properties, Phys.
+Rev.
+Lett.
+120, 145301 (2018).
+- Ye et al. [2024b] C.-Y. Ye, H.-M. Weng, and Q.-S. Wu, Con-cdvae: A method for the conditional generation of crystal structures, Computational Materials Today 1, 100003 (2024b).
+- Franceschetti and Zunger [1999] A. Franceschetti and A. Zunger, The inverse band-structure problem of finding an atomic configuration with given electronic properties, Nature 402, 60 (1999).
+- Cheng et al.
+[2023] G.
+Cheng, X.-G.
+Gong, and W.-J.
+Yin, Global optimization in the discrete and variable-dimension conformational space: The case of crystal with the strongest atomic cohesion, (2023), arXiv:2302.13537 [cond-mat.mtrl-sci] .
+- Note [2] For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling.
+- Zhao et al.
+[2017] X.-G.
+Zhao, J.-H.
+Yang, Y.
+Fu, D.
+Yang, Q.
+Xu, L.
+Yu, S.-H.
+Wei, and L.
+Zhang, Design of lead-free inorganic halide perovskites for solar cells via cation-transmutation, Journal of the American Chemical Society 139, 2630 (2017).
 - Ren et al.
@@
 - Köhler et al. [2023] J. Köhler, M. Invernizzi, P. De Haan, and F. Noé, Rigid body flows for sampling molecular crystal structures, (2023), arXiv:2301.11355 [cs.LG] .
+- Miller et al. [2024] B. K. Miller, R. T. Chen, A. Sriram, and B. M. Wood, Flowmm: Generating materials with riemannian flow matching, (2024), arXiv:2406.04713 [cs.CL] .
 - Gebauer et al. [2018] N. W. Gebauer, M. Gastegger, and K. T. Schütt, Generating equilibrium molecules with deep neural networks, (2018), arXiv:1810.11347 [stat.ML] .
+- Gebauer et al.
+[2019] N.
+Gebauer, M.
+Gastegger, and K.
+Schütt, Symmetry-adapted generation of 3d point sets for the targeted discovery of molecules, in Advances in Neural Information Processing Systems, Vol.
+32 (Curran Associates, Inc., 2019).
 - Gebauer et al.
@@
 Narasimhan, Tree of thoughts: Deliberate problem solving with large language models, (2023), arXiv:2305.10601 [cs.CL] .
-- Jain and Bligaard [2018] A. Jain and T. Bligaard, Atomic-position independent descriptor for machine learning of material properties, Phys. Rev. B 98, 214112 (2018).
 - Goodall et al.
@@
 Ono, Crystalformer: Infinitely connected attention for periodic structure encoding, (2024), arXiv:2403.11686 [cs.LG] .
+- Xu et al. [2024] H. Xu, D. Qian, and J. Wang, Predicting many properties of crystals by a single deep learning model, (2024), arXiv:2405.18944 [cond-mat.mtrl-sci] .
 - Rubungo et al.
@@
 Dieng, Llm-prop: Predicting physical and electronic properties of crystalline solids from their text descriptions, (2023), arXiv:2310.14029 [cs.CL] .
-- Maddox [1988] J. Maddox, Crystals from first principles, Nature 335, 201 (1988).
+- Groom et al. [2016] C. R. Groom, I. J. Bruno, M. P. Lightfoot, and S. C. Ward, The cambridge structural database, Structural Science 72, 171 (2016).
 - Brown et al.
@@
 Gadepally, Neural scaling of deep chemical models, Nature Machine Intelligence 5, 1297 (2023).
-- Stokes and Hatch [1989] H. T. Stokes and D. M. Hatch, Isotropy Subgroups of the 230 Crystallographic Space Groups (WORLD SCIENTIFIC, 1989).
 - Müller [2013] U.
 Müller, Symmetry Relationships between Crystal Structures: Applications of Crystallographic Group Theory in Crystal Chemistry, International Union of Crystallography Texts on Crystallography (OUP Oxford, 2013).
+- Stokes and Hatch [1989] H. T. Stokes and D. M. Hatch, Isotropy Subgroups of the 230 Crystallographic Space Groups (WORLD SCIENTIFIC, 1989).
 - Cox and White [2022] S. Cox and A. D. White, Symmetric molecular dynamics, Journal of Chemical Theory and Computation 18, 4077 (2022).
+- Dathathri et al.
+[2019] S.
+Dathathri, A.
+Madotto, J.
+Lan, J.
+Hung, E.
+Frank, P.
+Molino, J.
+Yosinski, and R.
+Liu, Plug and play language models: A simple approach to controlled text generation, (2019), arXiv:1912.02164 [cs.CL] .
+- Zhang et al.
+[2020] M.
+Zhang, N.
+Jiang, L.
+Li, and Y.
+Xue, Language generation via combinatorial constraint satisfaction: A tree search enhanced monte-carlo approach, (2020), arXiv:2011.12334 [cs.LG] .
+- Qin et al.
+[2022] L.
+Qin, S.
+Welleck, D.
+Khashabi, and Y.
+Choi, Cold decoding: Energy-based constrained text generation with langevin dynamics, Advances in Neural Information Processing Systems 35, 9538 (2022).
+- Lew et al. [2023] A. K. Lew, T. Zhi-Xuan, G. Grand, and V. K. Mansinghka, Sequential monte carlo steering of large language models using probabilistic programs, (2023), arXiv:2306.03081 [cs.AI] .
 - Lai et al.
@@
 Ke, End-to-end crystal structure prediction from powder x-ray diffraction, (2024), arXiv:2401.03862 [physics.chem-ph] .
-- Wang et al. [2021b] H.-C. Wang, S. Botti, and M. A. L. Marques, Predicting stable crystalline compounds using chemical similarity, npj Computational Materials 7, 12 (2021b).
-- Schmidt et al.
-[2022a] J.
-Schmidt, N.
-Hoffmann, H.-C.
-Wang, P.
-Borlido, P.
+- Raymond [1999] E. S. Raymond, The Cathedral and the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary (O’Reilly Media, Sebastopol, CA, 1999).
+- Court et al.
+[2020] C.
 J.
+Court, B.
+Yildirim, A.
+Jain, and J.
 M.
-A.
-Carriço, T.
-F.
+Cole, 3-d inorganic crystal structure generation and property prediction via representation learning, Journal of Chemical Information and Modeling 60, 4518 (2020), pMID: 32866381.
+- Davies et al.
+[2019] D.
+W.
+Davies, K.
 T.
-Cerqueira, S.
-Botti, and M.
-A.
-L.
-Marques, Large-scale machine-learning-assisted exploration of the whole materials space (2022a), arXiv:2210.00579 [cond-mat.mtrl-sci] .
-- Schmidt et al.
-[2022b] J.
-Schmidt, H.-C.
-Wang, T.
-F.
-T.
-Cerqueira, S.
-Botti, and M.
-A.
-L.
-Marques, A dataset of 175k stable and metastable materials calculated with the pbesol and scan functionals, Scientific Data 9, 64 (2022b).
+Butler, A.
+J.
+Jackson, J.
+M.
+Skelton, K.
+Morita, and A.
+Walsh, Smact: Semiconducting materials by analogy and chemical theory, Journal of Open Source Software 4, 1361 (2019).
+- [123] https://github.com/sparks-baird/matbench-genmetrics.
+- Goodfellow et al. [2016] I. Goodfellow, Y. Bengio, and A. Courville, Deep Learning (MIT Press, 2016) http://www.deeplearningbook.org.
 - Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized Gradient Approximation Made Simple, Phys. Rev. Lett. 77, 3865 (1996).
@@
 - Kresse and Furthmüller [1996] G. Kresse and J. Furthmüller, Efficient iterative schemes for ab initio total-energy calculations using a plane-wave basis set, Phys. Rev. B 54, 11169 (1996).
+- Chen et al.
+[2020] M.
+Chen, A.
+Radford, R.
+Child, J.
+Wu, H.
+Jun, D.
+Luan, and I.
+Sutskever, Generative pretraining from pixels, in Proceedings of the 37th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol.
+119, edited by H.
+D.
+III and A.
+Singh (PMLR, 2020) pp.
+1691–1703.
 Appendix A More details of CrystalFormer
 To recap, the space group information plays a key role in the architecture, training, and sampling of CrystalFormer.
-First of all, the vector embedding of space group number controls all subsequent outputs in the transformer that determines the Wyckoff letters, chemical species, fractional coordinates, and lattice parameters.
+First of all, the vector embedding of space group number controls all subsequent outputs of the transformer corresponding to the Wyckoff letters, chemical species, fractional coordinates, and lattice parameters.
 Second, the information of the space group and Wyckoff letter are used to select active components in the fractional coordinates and lattice parameters in the loss function during training.
-Lastly, the space group determines the concrete meaning of in terms of multiplicities and fractional coordinates.
-This information is used to place right amount of atoms to appropriate locations during sampling.
+Lastly, the space group determines the concrete meaning of Wyckoff letters in terms of multiplicities and fractional coordinates, which are used to place the right number of atoms precisely in the unit cell during sampling.
 A.1 Model architectures
@@
 A.2 Sampling algorithm
-Algorithm 2 summarizes the sampling method of CrystalFormer.
-It takes 520 seconds to generate a batch size 13,000 crystal samples on a single A100 GPU, which translates to a generation speed 40 millisecond per sample.
-Appendix B Stable crystal samples
-Table S2 list generated samples in the space group (No.
-225) with .
-Three of them (InNd, InTb, and EuTl) actually fall into the space group (No.
-221) due to merging of ”4a” and ”4b” Wyckoff positions that are occupied by the same elements.
-Overall, out of 1000 samples generated by CrystalFormer, we found 10 stable materials that are not contained in the Materials Project database.
-Extending the search to more recent material databases, we found match of AcCuRh, AcHgIn, AcAgSi and AuInPm in the WBM dataset [93], AcHgIn, AcAgSi, InPmTl, PdPmZn, IrLuPm and AuInPm in the Alexandria dataset [94, 95].
-Neverthless, we did not find AcBiK, InTb, and CaPaRu to be present in any of these dataset.
-To estimate the energy above hull, the DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [96] and all-electron projector-augmented wave method [97], as implemented in the VASP code [98].
+Algorithm 2 summarizes the autoregressive sampling method of CrystalFormer.
+It takes 520 seconds to generate a batch size 13,000 crystal samples on a single A100 GPU, which translates to a generation speed of 40 milliseconds per sample.
+Algorithm 3 summarizes the Markov chain Monte Carlo sampling of CrystalFormer.
+It is used in the structure constrained generation of crystals in Sec.
+IV.2.
+The property-guided materials design discussed in Sec.
+IV.3 employs a similar sampling strategy for the posterior probability distribution Eq.
+(3).
+Appendix B Validity and novelty of generated samples
+Figure S2 illustrates the structure and compositional validity of generated samples across all 230 space groups.
+Following the Ref.
+[121], a structure meets the validity criteria if the shortest atomic distance exceeds , a lenient standard.
+Composition validity requires charge neutrality as computed by SMACT [122].
+This is, however, an overly stringent criterion since the composition validity of the training set is only around by this measure [11].
+Note that the CrystalFormer can generate reasonable samples even for those space groups without any training data.
+This because the model can exploit knowledge learned from other space groups to place suitable atoms in the Wyckoff positions due to weight sharing.
+Moreover, since the sampling process makes use the of Wyckoff position table.
+The three dimensional coordinates of atoms are not completely random even for unseen space groups.
+| Space group | Crystal system | Training samples | Validity () | |
+| Struc. | Comp. | |||
+| 2 | Triclinic | 676 | 83.10 | 83.0 |
+| 12 | Monoclinic | 1273 | 87.70 | 81.80 |
+| 62 | Orthorhombic | 1187 | 95.50 | 87.20 |
+| 139 | Tetragonal | 1233 | 97.70 | 83.40 |
+| 166 | Trigonal | 1076 | 98.50 | 85.0 |
+| 194 | Hexagonal | 1129 | 99.40 | 89.90 |
+| 225 | Cubic | 3960 | 99.60 | 93.50 |
+| 1 | Triclinic | 27136 | 91.40 | 80.20 |
+| Autoregressive models | ||||
+| PGSchNet [100] | 99.65 | 75.96 | ||
+| LM-CH (character-level tokenization) [18] | 84.81 | 83.55 | ||
+| LM-AC (atom coordinate-level tokenization) [18] | 95.81 | 88.87 | ||
+| Crystal-text-LLM [20] | 96.5 | 86.3 | ||
+| Diffusion models | ||||
+| CDVAE [11] | 100.0 | 86.70 | ||
+| DiffCSP [13] | 100.0 | 83.25 | ||
+| DiffCSP++ [26] | 99.94 | 85.12 | ||
+| UniMat-Large [15] | 97.2 | 89.4 | ||
+Table LABEL:table:validity_table reports the validity of generated samples for selected space groups in each of the seven crystal systems.
+To ensure the numbers are representative, we chose the space group to be the one with the most training data for each crystal system.
+One sees that the model performs better for more symmetric space groups.
+This is a nice feature that complements existing crystal generative models, which mostly have difficulties in generating highly symmetric structures.
+As a reference, we also list the validity of the generated samples suppose one treats the crystals as if they are all in the space group (No.
+1) with only translational symmetry.
+One sees the structure validity scores of space group improve compared to the one shown in Figure S2 due to increased training samples.
+The second part of Table LABEL:table:validity_table shows reference results in the literature for the same validity test.
+In principle, the performance of the present model should fall back to the language model approaches [18, 20].
+The remaining gap may be due to details such as including the header line in the crystallographic information files (CIF), specific sampling strategy of language models, or the additional post-selection of samples [20].
+In the table, the DiffCSP++ [26] is the only alternative model that exploits the space group symmetry in a rigorous manner similar to ours.
+However, the DiffCSP++ does not predict Wyckoff positions in the generation process.
+Instead, one needs to search for template structures in the training set for generation, which may limit its generality.
+Besides works listed in Table LABEL:table:validity_table that reported validity scores in comparable settings, Ref.
+[19] has conditioned the generation of CIF on the space group symbols in a language model setting.
+Ref.
+[16] considered space group conditioned crystal generation using a fine-tuned generative model with space group labels.
+Neither approach provides exact constraints on the space group, which could yield problematic structures for large systems and highly symmetric space groups.
+Ref.
+[23] considered generating symmetric crystals in their Wyckoff representation.
+However, the model does not consider fractional coordinates and lattice parameters, so it requires a subsequent computational search to completely determine the crystal structure.
+Figure S1(a) shows the validity of generated samples as a function of sampling temperature.
+One clearly sees that reduced temperature increases the validity of samples at the cost of reducing the diversity [20].
+Figure S1(b) shows the novelty and uniqueness evaluated on generated samples with temperature.
+Novelty quantifies the proportion of new structures in the generated samples that were unseen in the training dataset.
+Uniqueness represents the percentage of distinct, non-redundant structures among the generated samples [123].
+One sees that across different temperatures the uniqueness remains high, indicating the model does not collapse to a mode that produces duplicated samples.
+On the other hand, the model produces close to novel material as temperature increases, which nicely demonstrates modal covering behavior of the maximum likelihood estimation training [124].
+Having a model distribution broader than the span of the dataset is crucial for material discovery.
+Appendix C Discovered crystal samples with symmetric structure initialization
+Table LABEL:table:samples and Table LABEL:table:pyxtal_samples list novel samples with eV/atom with CrystalFormer and PyXtal initializations respectively.
+Among them, Pd3Pt, MgAgPd2, TaNbRu2, Gd2HgAu, and TbMgPd2 are both discovered by CrystalFormer and PyXtal initialization.
+These compounds share identical crystal structures in the two tables, with lattice constants differing by less than 0.02 .
+To relax samples and estimate the energy above the hull, the DFT calculations were performed with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [125] and all-electron projector-augmented wave method [126], as implemented in the VASP code [127].
 All parameters of the calculations including settings of PBE functional, Hubbard U corrections, and ferromagnetic initialization are chosen to be consistent with Materials Project by using of MPRelaxSet function in pymatgen [81] A double relaxation strategy was employed.
 The maximum optimization ionic step and the maximum running time were constrained to 150 steps and 20 hours, respectively.
-All structures containing Yb element are ignored when calculating energy above hull due to they are unavailable from the Materials Project at the time of writing.
-| In MP | Not in MP | ||||
-| Formula | (eV/atom) | Lattice constant () | Formula | (eV/atom) | Lattice constant () |
-| PdPtTm | -0.00164 | 6.90 (6.86) | AcCuRh | -0.14248 | 7.47 |
-| LiPtZr | -0.01668 | 6.43 (6.39) | AcAgSi | -0.04364 | 7.81 |
-| CdLiPb | -0.00313 | 6.82 (6.79) | IrLuPm | -0.0289 | 6.89 |
-| LiNdPb | -0.05703 | 7.05 (7.03) | InPmTl | -0.01318 | 7.73 |
-| BeOsSi | -0.00779 | 5.77 (5.74) | PdPmZn | -0.0843 | 7.27 |
-| CeAs | -0.0073 | 5.99 (6.09) | AuInPm | -0.0568 | 7.70 |
-| Xe | -0.00067 | 7.45 (6.66) | AcBiK | -0.20215 | 8.37 |
-| InLiY | -0.17504 | 7.54 | InTb | -0.26278 | 7.95 |
-| InNd | -0.21393 | 8.35 | AcHgIn | -0.40448 | 8.66 |
-| BeC | -0.01227 | 4.33 (4.32) | CaPaRu | -0.09344 | 6.80 |
-| EuTl | -0.1054 | 7.90 | |||
+All structures containing Yb element are ignored when calculating energy above the hull due to they are unavailable from the Materials Project at the time of writing.
+| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) |
+| LuSiNi | 62 | c-Lu-c-Si-c-Ni | 0.0134 |
+| Eu3Mg | 62 | c-Eu-c-Mg-d-Eu | 0.0661 |
+| EuAgAu | 62 | c-Eu-c-Ag-c-Au | 0.0111 |
+| SmSbPd | 62 | c-Sm-c-Sb-c-Pd | 0.0368 |
+| Nd3Rh | 62 | c-Nd-c-Rh-d-Nd | 0.0673 |
+| SrSnPd | 62 | c-Sr-c-Sn-c-Pd | -0.0531 |
+| SrCdSn | 62 | c-Sr-c-Cd-c-Sn | 0.0863 |
+| SrMgHg | 62 | c-Sr-c-Mg-c-Hg | 0.0517 |
+| Eu2Ru | 62 | c-Eu-c-Eu-c-Ru | 0.0890 |
+| PrHgAu2 | 62 | c-Pr-c-Hg-c-Au-c-Au | 0.0868 |
+| Ce(SiPt)2 | 139 | a-Ce-d-Si-e-Pt | 0.0648 |
+| Dy(AlGa)2 | 139 | a-Dy-d-Al-e-Ga | 0.0902 |
+| Nd3Dy | 139 | a-Dy-b-Nd-d-Nd | 0.0690 |
+| Pd3Pt | 139 | a-Pt-b-Pd-d-Pd | -0.0298 |
+| LaMgSn | 139 | c-Mg-e-La-e-Sn | 0.0274 |
+| Rb3Mn2F7 | 139 | a-F-b-Rb-e-Rb-e-Mn-e-F-g-F | 0.0874 |
+| AlAgAu2 | 139 | a-Al-b-Ag-d-Au | 0.0912 |
+| Ba3Tl | 139 | a-Tl-b-Ba-d-Ba | 0.0933 |
+| CsPrTe2 | 166 | a-Cs-b-Pr-c-Te | 0.0977 |
+| Sr2Ca | 166 | a-Ca-c-Sr | 0.0771 |
+| Pr2AlRu3 | 166 | a-Al-c-Pr-d-Ru | 0.0909 |
+| Dy3Ho | 194 | d-Ho-h-Dy | 0.0738 |
+| Zr3Pb | 194 | d-Pb-h-Zr | 0.0426 |
+| PrAu3 | 194 | d-Pr-h-Au | 0.0696 |
+| Pd3Pt | 194 | d-Pt-h-Pd | 0.0053 |
+| TbZnGa | 194 | a-Tb-c-Ga-d-Zn | 0.0718 |
+| BaSr(GaGe)2 | 194 | a-Ba-b-Sr-f-Ga-f-Ge | 0.0855 |
+| Ho2Er | 194 | b-Er-f-Ho | 0.0790 |
+| NdDyHg2 | 225 | a-Nd-b-Dy-c-Hg | 0.0974 |
+| MgAgPd2 | 225 | a-Mg-b-Ag-c-Pd | 0.0540 |
+| PrEuIn2 | 225 | a-Eu-b-Pr-c-In | 0.0957 |
+| TaNbRu2 | 225 | a-Ta-b-Nb-c-Ru | 0.0823 |
+| PdAu | 221 | a-Au-b-Pd | 0.0025 |
+| Gd2HgAu | 225 | a-Hg-b-Au-c-Gd | 0.0522 |
+1
+The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/1gOIkWkjSH_Ed0-wzPk8VgxLjJcidbrH3/view?usp=sharing
+| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) |
+| Ba2Si | 62 | b-Ba-c-Ba-c-Si | 0.0766 |
+| GdIr | 62 | c-Gd-c-Ir | -0.2310 |
+| Nd3Dy | 139 | a-Dy-b-Nd-d-Nd | 0.0680 |
+| Pd3Pt | 139 | a-Pd-b-Pt-d-Pd | 0.0147 |
+| PdPt | 221 | a-Pd-b-Pt | 0.0847 |
+| NdDyHg2 | 225 | a-Dy-b-Nd-c-Hg | 0.0973 |
+| MgAgPd2 | 225 | a-Mg-b-Ag-c-Pd | 0.0480 |
+| HfRe | 221 | a-Re-b-Hf | 0.0926 |
+| TaNbRu2 | 225 | a-Nb-b-Ta-c-Ru | 0.0770 |
+| PdAu | 221 | a-Au-b-Pd | 0.0056 |
+| Gd2HgAu | 225 | a-Hg-b-Au-c-Gd | 0.0476 |
+| TbMgPd2 | 225 | a-Mg-b-Tb-c-Pd | 0.0676 |
+1
+The fractional coordinates and lattice parameters are omitted for brevity. See details at https://drive.google.com/file/d/18sdsp-6yRSBaNez1A0lr20Ru6H7OOzvJ/view?usp=drive_link.
+Appendix D Discovered crystal samples with element substitution
+Tables LABEL:table:mutation_samples and LABEL:table:mutation_pmg_samples list double perovskites crystals in the (No.
+225) space group with eV/atom found by CrystalFormer and element substitution of pymatgen respectively.
+We found that CrystalFormer has found three times more perovskites crystals compared to ionic substitution [46].
+The DFT calculation setup is the same as in the previous section.
+| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) | Lattice constant () |
+| Sr2TaNiO6 | 225 | a-Ni-b-Ta-c-Sr-e-O | 0.0426 | 7.9762 |
+| Ba2TaNiO6 | 225 | a-Ni-b-Ta-c-Ba-e-O | 0.0711 | 8.2129 |
+| Ba2YMnO6 | 225 | a-Y-b-Mn-c-Ba-e-O | 0.0284 | 8.3575 |
+| Eu2CoMoO6 | 225 | a-Co-b-Mo-c-Eu-e-O | 0.0720 | 7.9220 |
+| Ba2TbFeO6 | 225 | a-Tb-b-Fe-c-Ba-e-O | 0.0343 | 8.3200 |
+| Ba2HfWO6 | 225 | a-W-b-Hf-c-Ba-e-O | 0.0939 | 8.3568 |
+| Eu2CoNiO6 | 225 | a-Ni-b-Co-c-Eu-e-O | 0.0646 | 7.6181 |
+| Sr2NbVO6 | 225 | a-V-b-Nb-c-Sr-e-O | 0.0166 | 8.0375 |
+| Sr2NaWO6 | 225 | a-Na-b-W-c-Sr-e-O | 0.0901 | 8.2959 |
+1
+The fractional coordinates are omitted for brevity.
+| Formula | Space group | Wyckoff-Atom sequence 1 | (eV/atom) | Lattice constant () |
+| Ba2FeBiO6 | 225 | a-Bi-b-Fe-c-Ba-e-O | -0.0030 | 8.3808 |
+| Ba2NpCoO6 | 225 | a-Np-b-Co-c-Ba-e-O | 0.0241 | 8.3983 |
+| Ba2BiMoO6 | 225 | a-Bi-b-Mo-c-Ba-e-O | 0.0706 | 8.6322 |
+1
+The fractional coordinates are omitted for brevity.
+Appendix E Details of plug-and-play materials design
+We build two independent predictive models of band gap and formation energy following the generative pretraining strategy [128].
+Generative pretrain of CrystalFormer allows the model to extract crystal representations.
+Then, by attaching a regression neural network to its feature and training on property labels, the model can be used to predict the properties of crystals.
+For the regression task, the mean absolute error corresponds to a predictive model characterized by the Laplace distribution , where is a positive scale parameter.
+According to Bayes’ rule, the posterior probability for crystals given the property is:
+The crystal log-likelihood with a given property condition is then given by:
+It is instructive to see that property-guided samples are balanced between the CrystalFormer prior and the property prediction model likelihood.
+Here plays the role of guidance strength.
+For multi-property prediction, the likelihood function can be extended to:
+Note that MCMC sampling via the Metropolis algorithm does not involve the normalization factor of the conditional distribution.
+In our experiments, we set for band gap prediction and for formation energy prediction, corresponding to the inverses of the MAE values obtained from the regression model.